### v1

# XPUTimer: Anomaly Diagnostics for Divergent LLM Training in GPU Clusters of Thousand-Plus Scale

Abstract
The rapid proliferation of large language models has driven the need for efficient GPU training clusters. However, ensuring high-performance training in these clusters is challenging due to the complexity of software-hardware interactions and the frequent occurrence of training anomalies. Since existing diagnostic tools are narrowly tailored to specific issues, there are gaps in their ability to address anomalies spanning the entire training stack. In response, we introduce XPUTimer, a real-time diagnostic framework designed for distributed LLM training at scale. XPUTimer first integrates a lightweight tracing daemon to monitor key code segments with minimal overhead. Additionally, it features a diagnostic engine that employs novel intra-kernel tracing and holistic aggregated metrics to efficiently identify and resolve anomalies. Deployment of XPUTimer across 6,000 GPUs over eight months demonstrated significant improvements across the training stack, validating its effectiveness in real-world scenarios.
1 Introduction
The advent of large language models (LLMs) has revolutionized the deep learning training community, driving substantial advancements in artificial intelligence-generated content (AIGC). Recognizing their transformative potential to enhance user experiences, leading corporations are proactively leveraging LLMs to enhance a wide array of user-oriented services [1, 2, 3]. To meet the computational demands of LLM training, they construct large-scale training clusters comprising the latest GPUs interconnected via high-bandwidth links.
Figure 1 depicts the general training stack of the large-scale training cluster in modern corporations. As shown, the operations team manages low-level resources, and the infrastructure team delivers training optimizations[4, 5, 6, 7], with particular emphasis on parallel backbones[8, 9, 10]. Supported by these two teams, multiple algorithm teams focus on adapting LLMs for user-facing applications through various training methods [11, 12]. Notably, since LLMs do not solve all problems, the training cluster also supports other deep learning jobs, such as recommendation models and their specific parallel backbone, TorchRec [13].
Figure 1: The summarized training stack of large-scale training cluster in Ant Group, highlighting XPUTimer’s position.
Efficient distributed LLM training of large scale requires delicate collaboration among the teams across the stack. However, due to the involvement of thousands of software and hardware components, training anomalies frequently arise from various layers. These anomalies include obvious job failures and non-obvious training slowdowns, which should be resolved by different teams.
Specifically, the algorithm teams may unintentionally use incorrect training configurations, leading to training slowdowns. Meanwhile, the slowdowns may also come from unoptimized operators, which fall under the responsibility of the infrastructure team. Moreover, a significant number of job failures are caused by hardware errors. However, the current training stack lacks a dedicated diagnostic framework to address these anomalies. In its absence, resolving such issues requires cross-team investigations, resulting in high communication overhead and low resolution efficiency.
In order to design a deployable diagnostic framework in the training cluster, we identify three main challenges based on real-world operational experiences. C-1: Designing a long-running and lightweight diagnostic framework is challenging. Addressing non-obvious training slowdowns and low-level communication errors requires long-term and in-depth monitoring. However, exhaustive tracing mechanisms, such as PyTorch’s built-in profiler, impose a high runtime memory overhead, making it unsuitable for meeting long-running requirements. C-2: Detecting and diagnosing the underlying root cause is challenging. Errors or slowdowns in LLM training often manifest with similar symptoms, such as process hangs or decreased training speed, which obscure the actual problematic machine or code. C-3: Providing a backbone-extensible diagnostic mechanism is challenging. As shown in Figure 1, at least three backbones are commonly used across the LLM community, not to mention those supporting other deep learning models. Making diagnostic mechanisms extensible to multiple backbones is a significant challenge.
Faced with these challenges, previous efforts[14, 2, 3] targeting distributed LLM training fail to provide comprehensive solutions. This is because these approaches are narrowly designed for specific problems or scenarios, making them inadequate for addressing issues across the software-hardware stack. For example, FALCON [14] and C4D [3] are capable of diagnosing low-level network issues. However, our observations suggest that many slowdowns are caused by upper-level teams, as shown in Table 1. Likewise, tools such as MegaScale [2] are optimized for pre-training scenarios where teams collaborate closely and use only the Megatron backbone, limiting their adaptability to the advancing training stack.
To this end, we present XPUTimer, a real-time and holistic anomaly diagnostic framework designed for efficient distributed LLM training in GPU clusters of thousand-plus scale. XPUTimer consists of two components: a per-training-process tracing daemon and a diagnostic engine. To collect runtime data with low overhead (C-1), the tracing daemon instruments only key code segments rather than blindly tracing all runtime data. Furthermore, as the code segments are carefully selected and intercepted at the level of both Python and C++ runtime, the instrumentation provides sufficient critical information for diagnosis (C-2) while remaining backbone-agnostic (C-3). Notably, XPUTimer is also hardware-extensible, which could provide seamless support for other NPUs.
With the real-time data collected by the tracing daemon, the diagnostic engine detects and diagnoses anomalies, including errors and slowdowns(C-2). For error diagnostics, XPUTimer introduces a novel intra-kernel tracing mechanism, providing fine-grained diagnostics specifically targeting communication-related hang errors. Compared to a binary search using NCCL tests [5, 2], it reduces the complexity of faulty machine diagnostics from $O(\log N)$ to $O(1)$. For slowdown diagnostics, XPUTimer proposes holistic aggregated metrics that encompass not only commonly used macro metrics like training throughput but also novel micro metrics, such as issue latency distribution. With these metrics, XPUTimer can diagnose non-obvious slowdowns (e.g., $2.66\%$) in real-world workloads.
[…]
XPUTimer has been open-sourced at https://github.com/intelligent-machine-learning/dlrover/tree/master/xpu_timer. XPUTimer serves as a core component of DLRover [15], an automated distributed DL system deployed within Ant Group and supported by the LF AI & Data Foundation [16].
Our contributions are as follows.
  • We highlight the urgent need for a real-time, holistic diagnostic framework capable of identifying LLM training anomalies across the entire training stack.
  • We introduce XPUTimer, a diagnostic framework specifically designed to tackle the critical challenges of long-term monitoring, root cause diagnosis, and backbone extensibility in LLM training diagnostics.
  • We deploy XPUTimer across more than 6,000 GPUs over an 8-month period, deriving typical case studies and practical insights from its daily operations.
Table 1: A comprehensive analysis of common anomalies encountered in Ant Group, annotated with XPUTimer’s primary target.
Anomalies
Type Error Slowdown
Taxonomy
Algorithm
bugs
Infrastructure
bugs
OS
errors
GPU
errors
Network
errors
New
algorithms
Unnecessary
synchronization
Un-optimized
kernels
Memory
management
GPU
underlocking
Network
jitter
Symptom
Startup crash or
hang error
Runtime hang
or crash error
Slowdown compared to historical
jobs and prior training steps
Slowdown compared to
prior training steps
Team Algorithm Intrastructure Operations Algorithm Intrastructure Operations
Attribution Obvious Obscure Obvious
Comment No need to diagnose XPUTimer’s main target
2 Background and Motivation
In this section, we introduce the training of LLMs within corporations like Ant Group, and motivate the design of XPUTimer for anomaly diagnostics of LLM training.
2.1 Large-Scale LLM Training Stack
The pursuit of more powerful large language models [17, 18, 19, 20, 21, 22, 23, 24, 25, 26] has sparked an intense competition between leading corporations. Referring to the software-hardware stack for training LLMs in Figure 1, we delve into how LLMs are reshaping corresponding teams.
Advancing of LLM applications.
With the advancing of LLM algorithms [18], LLMs excel not just in natural language understanding, but also in tackling advanced tasks such as multimodal tasks  [27, 28], reasoning tasks [29, 30]. For instance, customer service teams fine-tune LLMs to develop chatbots that generate accurate responses to complicated questions. Similarly, product development teams leverage multimodal LLMs to generate comprehensive product descriptions by integrating diverse inputs, such as product images and textual data. Consequently, different algorithm teams continue to innovate LLM models [23, 24, 25, 26, 18, 31, 32, 33, 34, 35, 11, 36] tailored for diverse application scenarios.
Advancing of training cluster.
Training these LLMs is computationally intensive, requiring large-scale GPU clusters of thousand-plus scale [26]. Therefore, leading corporations are continuously investing in large training clusters powered by state-of-the-art hardware [37, 38]. These clusters not only expand in scale but also incorporate the latest GPUs [39, 40] with higher computational performance and advanced interconnect [41, 42]. Efficient operation of these large-scale clusters is critical to the corporations, which requires the operations team to ensure uninterrupted training performance. Its responsibilities include job scheduling, driver updates, hardware maintenance, etc.
Advancing of training infrastructure.
To enable easy, efficient, and scalable LLM training within large-scale clusters, the infrastructure teams build the dedicated software stack. This stack bridges the gap between algorithm teams and operations team. Specifically, the infrastructure team focuses on optimizing the software stack by integrating advanced operator libraries [4, 43, 5], training framework [7], and state-of-the-art model parallel backbones [8, 9, 44, 13]. By leveraging this highly optimized training infrastructure, we unlock the full potential of LLMs, enabling innovation and scalability across our various services.
2.2 Anomalies of Large-scale LLM Training
Due to the complexity of the entire training stack, various issues can easily occur. These issues affect both the training speed of individual training jobs and the overall utilization of training clusters, collectively termed as anomalies. Table 1 presents a distilled analysis of the common anomalies in our real-world cluster, broadly categorized into two primary types: errors and slowdowns. These anomalies often span responsibilities across multiple teams  [2, 3, 14], making them notoriously difficult to diagnose. Thus, resolving these anomalies demands a diagnostic framework that can effectively identify their attributions. However, designing such a framework is far from trivial, posing three key challenges.
Figure 2: Slowdown of a 1024-GPU training job with Llama2-70B under various single-GPU underclocking configurations. Results are evaluated using both Megatron [9] and FSDP [8].
The pain of triggering condition.
Anomalies in large-scale LLM training can arise at any stage and time, as shown in the fourth row of Table 1. First, low-level OS or hardware issues, often attributed to the operations team, may occur randomly, causing runtime hangs or slowdowns. For example, Figure 2 illustrates a significant MFU decline resulting from single-GPU underclocking. Second, training slowdowns caused by codebase modifications may also occur randomly. Specifically, codebase updates on the LLM algorithm typically cause slowdowns at the start, while modifications related to the training infrastructure can result in mid-training slowdowns. A notable example is the slowdown caused by on-demand garbage collection (GC) [2] of Python runtime, as detailed in §5.2.2. In such cases, a long-term background diagnostic mechanism is required to identify these anomalies. However, given the variety of anomalies across the stack, it is challenging to design a lightweight diagnostic framework capable of efficiently collecting key metrics.
The pain of obscured attribution.
[…]
The pain of backbone extensibility.
With decades of advancements in model parallelism for large-scale training, a variety of divergent backbones have been developed to parallelize LLMs. Most LLMs are pre-trained using Megatron [9], a specialized parallel backbone designed for transformer-based LLMs. However, researchers continue to investigate alternative backbones [8, 10, 44] for fine-tuning pre-trained LLMs on downstream tasks or for training more complex large models, such as multimodal models. For instance, fully sharded data parallelism (FSDP)[8] and DeepSpeed[10] are widely utilized within our cluster for training multimodal LLMs. These alternatives are often selected to better align with specific training requirements. However, integrating the diagnostic framework into divergent parallel backbones remains challenging, as their runtime environments and programming interfaces vary significantly.
Existing works have made efforts to build a more robust training stack for large-scale LLM training [2, 3]. They generally operate under several assumptions: (1) training tasks rely on limited backbones like Megatron, (2) all teams work in close collaboration to pre-train a single LLM, or (3) The focus is on specific anomalies, such as communication-related issues. These assumptions fail to reflect reality and these works cannot solve the problems across the full stack. To this end, we propose XPUTimer, a cluster-wide solution deployed in Ant Group’s training cluster. It monitors the real-time progress of LLM training jobs and diagnoses various anomalies precisely.
3 XPUTimer Design
Figure 3: Architecture overview of XPUTimer.
In this paper, we propose XPUTimer, a holistic and real-time diagnostic framework that identifies anomalies in large-scale training with low overhead and pinpoints their corresponding root causes. XPUTimer is designed and implemented following three principles.
  • XPUTimer should be lightweight enough. In this case, it could run as a long-term diagnostic mechanism without compromising the training performance at any scale.
  • XPUTimer should cover enough runtime information. In this case, it could identify the attribution precisely with detailed analysis when encountering anomalies.
  • XPUTimer should be backbone-agnostic. In this case, it is sufficiently extensible to function as a cluster-wide solution for anomaly diagnostics, regardless of the backbones employed.
Architecture overview.
Figure 3 illustrates the architecture of XPUTimer, which is deployed in Ant Group’s large-scale training cluster. XPUTimer is composed of two components: the tracing daemon and the diagnostic engine.
By automatically attaching a tracing daemon to each training process in LLM training jobs, XPUTimer enables a framework-agnostic and lightweight tracing mechanism. Specifically, the tracing daemon instruments only key code segments at the level of Python and C++ runtime, focusing on carefully selected Python APIs and GPU kernels. During instrumentation, events are injected to measure the latencies of these code segments. This approach draws on our experience hosting various large-scale LLM training jobs, targeting critical paths that commonly impact training efficiency.
Then, the timing data collected by the daemons is transmitted to the diagnostic engine for holistic anomaly diagnostics. The engine employs a fast hang-error diagnostic method and leverages novel aggregated metrics to effectively identify slowdowns. The aggregated metrics encompass both macro-level metrics, such as training throughput, and micro-level metrics, including issue latency distribution of GPU kernels. Once the diagnostic engine identifies errors and slowdowns, the issues are routed to attribution teams, enabling swift resolution.
4 Lightweight Selective Tracing
To collect sufficient real-time data for anomaly diagnosis in the LLM training cluster, XPUTimer’s tracing daemon offers backbone-agnostic and lightweight full-spectrum tracing. Its design focuses on two key aspects: determining what information to collect and establishing how to collect it efficiently.
4.1 Key Segment Instrumentation
[…]
Figure 4: Instrumented key code segments in XPUTimer.
Therefore, XPUTimer selectively instruments code segments of key APIs and kernels to collect real-time information. This design is based on an insight into LLM training on large-scale GPUs: LLM training is predominantly dominated by a limited set of deep learning operators. These operators mainly include matrix multiplication and cross-GPU communication operators. Figure 4 presents the specific code segments instrumented by XPUTimer for efficient anomaly diagnostics.
As shown in the figure, the instrumented code segments can be broadly categorized into two groups. The first category involves intercepting key API calls, including those related to Python’s garbage collection (GC), PyTorch’s dataloader, and GPU synchronization. These APIs are carefully selected based on empirical insights into performance issues and optimization opportunities. These insights are detailed in §5.2, with corresponding cases discussed in §7.
The second category focuses on intercepting critical GPU computation and communication kernels executed at the C++ runtime level. These kernels, primarily provided by optimized libraries[4, 6, 46, 5], account for the majority of the workload during large-scale training. Additionally, there are customized kernels developed by the infrastructure team.
Notably, the above design enables XPUTimer to support backbone-agnostic and extensible tracing capabilities. Extending tracing capabilities for Python-related APIs is straightforward, requiring only the configuration of the specific environment variables in the training scripts, as shown below. {minted}[ frame=none, obeytabs=true, framesep=0mm, baselinestretch=0.8, fontsize=, xleftmargin=1.6em, breaklines, escapeinside=||, ]shell export TRACED_PYTHON_API="torch.cuda@synchronize" Meanwhile, intercepting C++ kernels necessitates explicit registration through a C++ interface. This requirement is feasible, as the infrastructure team takes charge of the development of both these customized operators and XPUTimer, ensuring seamless integration and functionality.
4.2 Timing in the Background
With intercepted Python APIs and GPU kernels, FLARE measures their elapsed latencies, as shown in Figure 5. Specifically, a dedicated tracing thread runs in the background to efficiently manage timing data. It employs different timing mechanisms for Python APIs and GPU kernels.
For synchronous Python API calls, FLARE directly records their start and end timestamps and forwards them to the timing manager. For GPU kernels, which execute asynchronously, FLARE injects CUDA events[47] after an interception to record execution status. These events are enqueued for further processing. The timing manager queries the status of the queued events in the background, avoiding any disruption to the training thread. Additionally, during GPU kernel interception, FLARE extracts input specifications, such as memory layout, to support subsequent anomaly diagnostics in §5.2.1.
Figure 5: Intercepting and timing the training in the background.
As training progresses, the timing manager proactively transmits all real-time data to XPUTimer’s diagnosis engine. By employing key segment instrumentation and running timing tasks in the background, XPUTimer minimizes both computing resource and memory overhead, ensuring efficient data collection for real-time anomaly diagnostics. A detailed evaluation of XPUTimer’s real-time overhead is provided in §6.1.
5 Anomaly Detection and Diagnosis
Using the real-time data collected by the tracing daemon, XPUTimer’s diagnostic engine identifies and analyzes anomalies encountered during distributed LLM training. In this section, we present XPUTimer’s holistic diagnostic workflow for addressing two common anomaly symptoms: errors and slowdowns.
5.1 Fast Runtime Error Diagnosis
As shown in the left of Table 1, errors encountered at the beginning of a training job are typically caused by bugs in the training scripts, which can often be addressed easily by the algorithm teams and infrastructure team. However, diagnosing errors that occur during the training progress is more challenging and critical. Such errors often stem from issues like operating system crashes, GPU failures, or network disruptions, which can generally be resolved by isolating the problematic machines and restarting the training job.
A typical symptom associated with these errors is the hanging of the training job. Training LLMs across numerous GPUs in a distributed manner inherently relies on the coordination of training processes. When the aforementioned errors occur, they rarely affect all training processes simultaneously. In this context, XPUTimer focuses on rapidly diagnosing hang errors by identifying faulty machines. Then, XPUTimer routes this information to the operations team, enabling the training job to restart with healthy machines.
Specifically, XPUTimer’s diagnostic engine first detects hang errors by examining the status of tracing daemons. The tracing daemon operates in the background of the training thread and continuously queries events recorded during job execution. If it fails to confirm the completion of an event within a predefined timeout interval, it proactively reports a potential hang error to the diagnostic engine. Similarly, if a tracing daemon does not transmit any real-time data within the specified timeout interval, the diagnostic engine also interprets this as an indication of a hang error.
After hang errors are reported, they are classified as either communication or non-communication errors. XPUTimer diagnoses these errors in two steps: first, a coarse-grained diagnosis through call stack analysis; and second, a fine-grained diagnosis using intra-kernel tracing
Diagnosis using call stack analysis.
This diagnosis is used to identify problematic machines encountering non-communication errors. Figure 6 illustrates an example of hang-error diagnosis via call stack analysis. As shown in the left of Figure 6, when the training process of rank-00 crashes or is suspended due to a non-communication error, it halts at a call stack corresponding to a non-communication function. In contrast, the training processes of other ranks continue executing correctly and eventually stop at a call stack associated with a communication-related function that depends on coordination with rank-00. In this scenario, the machine associated with rank-00 is identified as the source of the error. It should be noted that, although these non-communication errors may cause direct crashes, the call stack analysis could also locate the faulty machine.
However, communication hang errors cannot be identified through call stack analysis. As shown in the right of Figure 6, the training processes of all ranks terminate at the same call stack corresponding to a communication function, such as allreduce or allgather. In this scenario, there are no distinct differences between ranks based on call stack analysis.
Figure 6: Diagnosing hang errors via call stack analysis.
We further investigate the symptoms of communication hang errors and obtain two observations. Firstly, some communication hang errors generate error logs. For instance, if the link between RDMA NICs breaks, an error code of 12 is produced. Secondly, more hang errors result in an endless loop within the launched communication kernels, ultimately leading to job termination after a predefined timeout. To identify the unhealthy machine responsible for such errors, a straightforward solution is to perform a binary search by executing communication tests across all involved GPUs. This approach has a complexity of $O(\log N)$ and requires hours to pinpoint the faulty machine among thousands of GPUs [14].
Diagnosis using intra-kernel tracing.
Faced with this problem, XPUTimer introduces a minute-level diagnostic approach using intra-kernel tracing. This intra-kernel tracing leverages CUDA-GDB, the debugging tool for CUDA programming.
Specifically, XPUTimer’s diagnostic engine instructs the tracing daemon to attach the halted training processes with CUDA-GDB before terminating them. Once attached, the tracing daemon executes a script capable of automatically extracting detailed communication statuses to identify unhealthy machines. Figure 7 depicts an example of diagnosing communication hang errors in a hanging ring-allreduce kernel.
In the ring-allreduce kernel, each thread block is responsible for transmitting data between linked adjacent ranks within the kernel’s constructed ring. The data are split into chunks and thread blocks of adjacent ranks work together to transmit the chunks step by step. Thus, XPUTimer could retrieve the register values corresponding to the loop steps used for data transmission between linked ranks. Theoretically, the connection with the minimum step reveals the related GPUs experiencing errors. This intra-kernel tracing process is performed in parallel across all involved GPUs. As a result, its complexity is $O(1)$, enabling completion within a few minutes.
XPUTimer then routes the diagnostic information for detected errors to the operations team, assisting with tasks such as isolating faulty machines and restarting the training job.
5.2 Aggregation for Slowdown Diagnosis
As demonstrated in §2.2, slowdowns can be attributed to changes across the entire training stack. Meantime, slowdowns caused by software changes introduced by the algorithm and infrastructure teams are often subtle and challenging to detect. Identifying these changes typically requires comparisons across historical training jobs and prior training steps. In contrast, hardware changes, such as GPU underclocking or network jitter, are more apparent and can be detected solely through comparisons across training steps.
To holistically identify these anomalies, XPUTimer aggregates real-time data collected from the tracing daemon into five primary metrics, shown in Figure 8. These metrics are based on the consensus that a “healthy” training pipeline should exhibit a timeline saturated with GPU kernels dedicated to either computation or communication. Computation kernels should achieve high FLOPS, while communication kernels are expected to utilize high bandwidth. Any deviations from these characteristics point to idle GPU resources, signaling potential slowdowns in training jobs. Of the five metrics, three are commonly used in existing works [2, 14], while the other two are newly introduced by XPUTimer.
Figure 7: An example of diagnosing communication hang errors in a ring-allreduce kernel using intra-kernel tracing.
Figure 8: A timeline of a distributed training job annotated with aggregated metrics used for diagnosing slowdowns in XPUTimer.
5.2.1 Diagnosing Obvious Slowdown
1 Training throughput for detecting slowdown.
Training throughput is the most straightforward metric for detecting slowdowns. XPUTimer measures training throughput by timing the rate at which input data is consumed by the training pipeline. This is achieved by instrumenting the dataloader API of Pytorch. As a macro performance metric, training throughput directly reflects slowdowns in training efficiency through comparison to historical training jobs and between training steps of the same job. However, XPUTimer cannot diagnose the specific factors contributing to the slowdown. To address this, XPUTimer relies on the following four micro metrics to further investigate the underlying causes.
2 FLOPS for slow critical kernel.
XPUTimer monitors the FLOPS of instrumented critical computation kernels, leveraging timing data and input layout. By comparing the FLOPS of identical kernels across different ranks, XPUTimer diagnoses GPUs that exhibit poor computational performance, often caused by issues like GPU underclocking. Machines affected by GPU underclocking are then routed to the operations team for isolation. Additionally, by analyzing FLOPS, XPUTimer identifies un-optimized kernels in training jobs, particularly those with large input sizes but low FLOPS. These anomalies are detected without interrupting training jobs and are subsequently routed to the infrastructure team for further investigation. Notably, when analyzing FLOPS data, XPUTimer accounts for the impact of communication kernels that overlap with computation kernels.This ensures that computation kernels with falsely low FLOPS are not mistakenly flagged.
3 Bandwidth for slow connection.
XPUTimer monitors the bandwidth of communication kernels. A communication operator requires launching the communication kernels on all ranks. Since variations in kernel-issue timestamps exist across different ranks, XPUTimer calculates the communication bandwidth by utilizing the start and end timestamps of the final communication kernels issued across all participating ranks. The captured communication bandwidth is compared with offline profiled data. If low-bandwidth communication is detected, XPUTimer conducts a communication test using binary search to pinpoint machines experiencing issues such as network congestion. These slowdowns are then identified and routed to the operations team for resolution.
5.2.2 Diagnosing Obscured Slowdown
While the above three metrics ensure that both critical computation and communication GPU kernels operate at high performance, they do not cover the less critical operations, such as various CPU operations and element-wise activation GPU kernels. Meantime, XPUTimer’s selective key segment instrumentation also omits the monitoring of these operations.
To diagnose their potential contributions to slowdowns, we further classify these not-instrumented operations into three categories: intra-step CPU operations, inter-step CPU operations, and minority GPU kernels. Intra-step CPU operations and inter-step CPU operations differ due to their occurrences within the timeline of training steps. Minority GPU kernels refer to those GPU kernels that often occupy little GPU computation resources. Specifically, two metrics are introduced for the diagnostics: issue latency distribution for intra-step CPU operations and void percentage for inter-step CPU operations and minority GPU kernels.
Figure 9: Runtime overhead in terms of latency with various models, backbones, and number of GPUs.
4 Issue latency distribution for kernel-issue stall.
In a well-optimized parallel backbone, only the necessary intra-step CPU operations for launching GPU kernels or coordinating the training processes are expected. However, algorithm teams may inadvertently introduce unnecessary GPU synchronizations when modifying the LLM model. Meantime, certain function calls, such as GC [2, 9], may be implicitly triggered by the Python runtime. These intra-step CPU operations can occur repeatedly during the model’s forward pass, bringing considerable overhead. In such cases, these operations cause an anomaly known as a kernel-issue stall, leading to GPU idle time within the training step.
4
–1 in Figure 8 shows the example of Python runtime GC. In the figure, the Python runtime GC stalls the CPU thread and causes the lagging of GPU kernels on rank-1. Although the communication kernel on rank-0 is issued without stalling, it simply waits for the one on rank-1, ultimately causing the overall training speed to decline. 4–2 in Figure 8 shows an example of unnecessary GPU synchronization introduced by the developers from the algorithm teams. As all ranks wait for the completion of communication kernels, the kernel issue of follow-up kernels is stalled and not overlapped with GPU computation. When such unnecessary synchronization occurs repeatedly across the model’s forward pass, it ultimately results in a slowdown of the training speed.
Originally, detecting these anomalies of kernel-issue stall requires investigating the aggregated timeline with much human effort. Faced with this issue, XPUTimer proposes a new metric, named issue latency distribution, for diagnosing this issue without human intervention. Kernel-issue latency is defined as the time elapsed between the kernel’s issue timestamp and the start timestamp of its execution on the GPU. Based on our observation of anomalies of kernel-issue stall, the kernel-issue latencies of unhealthy training jobs should be much shorter than those of a healthy training job.
By monitoring runtime issue latency distribution, and comparing it with the historical data, XPUTimer could identify the anomaly of kernel-issue stall. Then, XPUTimer routes them to algorithm and infrastructure teams for resolution, as they are commonly software issues. §6.3 and §7.2 demonstrate the effectiveness of issue latency distribution in diagnosing kernel-issue stalls, even when the slowdown is minimal.
5 Void percentage for other un-covered operations.
While the tracing daemon only instruments the critical operators, inter-step CPU operations and minority GPU kernels both manifest as empty time slots in the visualized timeline, as shown in Figure 8. Consequently, XPUTimer introduces a metric, termed the void percentage, to identify slowdowns caused by these factors.
As for inter-step CPU operations, as depicted by 5–2 in Figure 8, XPUTimer measures the latency between the last kernel preceding the dataloader and the first kernel following the same dataloader. XPUTimer then computes the void percentage for inter-step CPU operations using the following equation:
$V_{inter}=T_{inter}\ /\ T_{step}$ (1)
where $T_{inter}$ represents the latency associated with inter-step CPU operations, and $T_{step}$ denotes the total latency of the training step.
As for minority GPU kernels, as shown by 5–1 in Figure 8, XPUTimer first automatically detects empty slots where GPU kernels are launched but remain un-executed. These empty slots signify that the GPUs are occupied by kernels outside the scope of XPUTimer’s tracing mechanism. XPUTimer subsequently accumulates these slots for each training step and computes the void percentage using the following equation:
$V_{minority}=T_{minority}\ /\ (T_{step}-T_{inter})$ (2)
where $T_{minority}$ is the latency of all minority GPU kernels.
When the void percentages ($V_{inter}$ and $V_{minority}$) surpass the predefined thresholds for a specific parallel backbone, XPUTimer annotates the training job with potential slowdowns attributed to inter-step CPU operations or minority GPU kernels. Then, XPUTimer notifies the algorithm and infrastructure team for further investigation.
6 Evaluation
In this section, we present experiments to demonstrate XPUTimer’s effectiveness from various perspectives. Specifically, we first evaluate XPUTimer’s runtime overhead in terms of latency and memory, showcasing its lightweight nature for runtime anomaly diagnosis. Then, we assess the effectiveness of XPUTimer’s novel diagnostic mechanisms: intra-kernel tracing for communication hang error diagnosis and issue distribution metric for kernel-issue stall diagnosis.
6.1 Runtime Overhead
We evaluate XPUTimer on four parallel backbones: Megatron [9], FSDP [8], DeepSpeed [10], and TorchRec [13]. Among these, Megatron, FSDP, and DeepSpeed are widely used for LLM training, while TorchRec is employed for training large recommendation models within Ant Group. Four models are benchmarked, spanning language, vision, and recommendation tasks: two large language models (Llama 18B and 70B), one large vision model (Llama Vision 40B), and one recommendation model (DLRM 72M).
The latency overhead experiment is conducted on 1,024 H800 GPUs deployed across 128 servers with RoCE connectivity. Figure 9 presents the latency overhead introduced by XPUTimer with various models, backbones, and number of GPUs. As shown, XPUTimer incurs a latency overhead of $0.43\%$ for three LLM training backbones and $1.02\%$ for TorchRec.
The memory overhead experiment is conducted on two setups, which are 16 A100 GPUs on 2 nodes and 1536 H800 GPUs on 192 nodes. We compare XPUTimer with Torch Full, Torch w/o Stack, Torch w/o Layout&Stack, and XPUTimer. Torch w/o Stack refers to using the PyTorch builtin profiler with stack tracing disabled, while Torch w/o Layout&Stack further disables matrix layout tracing.
[…]
From the above results, we can conclude that, regardless of the GPU scale, parallel backbone, parallel strategy, or model type used, XPUTimer consistently maintains an extremely low runtime overhead in terms of both latency and memory. The lightweight selective tracing facilitates XPUTimer’s deployment within our large training cluster, serving as a diagnostic framework for diverse training jobs.
Figure 10: Memory consumption of dumped logs per GPU per step using the PyTorch profiler and XPUTimer while training a Llama-70B model on 16 A100 GPUs.
6.2 Effectiveness of Intra-kernel Tracing
We evaluate the intra-kernel tracing mechanism on 16 A100 GPUs across two servers with RoCE connectivity. Given that most communication kernels are ring-based, this experiment focuses on evaluating ring-allreduce. Specifically, we customize the training script composed solely of communication kernels, with one GPU intentionally suspended to simulate a hang error caused by communication issues.
Figure 11 illustrates the pinpointing latencies for intra- and inter-server communication. The figure presents the latency results for three communication protocols [48] and cross-node configurations. As shown, XPUTimer requires $29.4\text{\textasciitilde}309.2s$ to detect erroneous GPUs across different scenarios. Among the protocols, XPUTimer performs best when the SIMPLE protocol is used for communication. This is because, with the SIMPLE protocol, XPUTimer only needs to scan the first thread of each thread block to check the steps, whereas the other two protocols require scanning the entire thread block.
When comparing intra-server and inter-server results, XPUTimer performs better when the ring-allreduce operation spans multiple servers. This is because intra-server GPUs are connected via NVLink, whereas inter-server GPUs communicate through NICs. Communication kernels launched over NICs involve fewer thread blocks, as NICs have fewer internal links compared to NVLink. As a result, XPUTimer scans fewer thread blocks for error diagnosis in inter-server scenarios.
In summary, the intra-kernel tracing mechanism can detect erroneous GPUs in a maximum of $309.2s$. Notably, as the complexity of intra-kernel tracing is $O(1)$, these results remain consistent regardless of scale.
Figure 11: Latency for pinpointing the erroneous GPUs causing a hang error in ring-allreduce with different protocols.
Figure 12: Issue distribution across all communication kernels for a Llama-20B model trained with Megatron and 256 GPUs, including the overall CDF and the CDFs for each type of communication kernel, respectively.
6.3 Effectiveness of Issue Latency Distribution
In this experiment, we evaluate the issue latency distribution using Llama-20B running on 256 H800 GPUs across 32 servers connected via RoCE. Figure 12 illustrates the issue latency distribution for all communication kernels in the Unhealthy-GC, Unhealthy-Sync, and Healthy scenarios. In the Unhealthy-GC scenario, GC is implicitly triggered by the Python runtime. In the Unhealthy-Sync scenario, an unintended GPU synchronization call is added within the transformer block, leading to repetitive GPU synchronizations during the model’s forward pass. In the Healthy scenario, GC is efficiently managed by the parallel backbone, and no unnecessary synchronizations are introduced.
As shown in the figure, the issue latency distribution patterns align with our claim in §5.2.2. The issue latency CDF of a healthy LLM training job increases linearly, whereas the issue latency CDFs for Unhealthy-GC and Unhealthy-Sync exhibit a much steeper rise. This is because the issue latencies of different ranks in the healthy scenario are solely influenced by the collective communication operator, resulting in a uniform distribution. In contrast, in the cases of Unhealthy-GC and Unhealthy-Sync, while some ranks are affected, their latencies become very short due to the delayed start of the issue time.
Since both GC and GPU synchronizations span the entire model forward pass, all communication kernels are affected, as illustrated in Figure 12. Furthermore, each training process triggers GC independently, and the GC operation for a single process is more time-consuming than GPU synchronization. Consequently, the issue latency distribution for Unhealthy-GC is worse than that of Unhealthy-Sync.
7 Deployment & Case Studies
In this section, we further demonstrate XPUTimer’s effectiveness in diagnosing LLM training anomalies through statistical analysis of its cluster-wide deployment and detailed case studies of various anomalies.
7.1 Cluster-wide Deployment
XPUTimer has been deployed within a training cluster with 6,000 GPUs for over 8 months. During this period, it is responsible for monitoring, detecting, and diagnosing training jobs for various deep learning models, especially large-scale distributed LLM training. Its diagnostic capabilities enable the algorithm teams, infrastructure team, and operations team to seamlessly enhance model training efficiency and improve training cluster utilization.
Table 2: Typical errors detected by XPUTimer.
Taxonomy Details Numbers Mechanism
OS errors Checkpoint storage 10 Stack analysis
OS crash 1
GPU errors GPU Driver 26
Faulty GPU (Unknown) 37
NCCL hang 36 Intra-kernel tracing
Network errors
RoCE issue 17
Table 2 presents a subset of error anomalies detected by XPUTimer. As shown in the table, XPUTimer effectively identifies OS- and hardware-related anomalies, including crashes, hangs, and slowdowns caused by such issues. While these issues are typically conspicuous and could be detected by existing methods based on noticeable training interruptions [9, 14], XPUTimer’s novelty lies in providing richer error information through techniques like intra-kernel tracing. Such runtime information helps ease and accelerate the attribution process of low-level issues for operations teams.
Moreover, the true value of XPUTimer lies in its ability to detect anomalies caused by both algorithm teams and the infrastructure team. Table 3 summarizes slowdowns diagnosed by XPUTimer using its aggregated metrics. In the table, anomalies newly identified by XPUTimer, as compared to prior works [2], are highlighted in bold. The following subsections provide a detailed discussion of typical cases.
Table 3: Slowdowns diagnosed by XPUTimer, with “Details” showing training job specifics and associated MFU decline.
Metric Attribution Details
FLOPS GPU underclocking 480 GPUs, Llama-65B, $14\%\downarrow$
Backbone migration 1856 GPUs, Llama-80B, $33.3\%\downarrow$
Bandwidth
Network jitter with
increased CRC
928GPUs, Llama-65B, $10\text{\textasciitilde}20\%\downarrow$
Down of
GDR module
32GPUs, Llama-10B, $80\%\downarrow$
128GPUs, Llama-10B, $62.5\%\downarrow$
Host-side hugepage
caused high sysload
128GPUs, LlamaVision-11B, $20\%\downarrow$
Issue latency distribution Python GC
2048GPUs, Llama-80B, $10\%\downarrow$
280GPUs, LlamaVision-11B, $60\%\downarrow$
Unnecessary
GPU Sync
256GPUs, Llama-20B, $2.66\%\downarrow$
Package chcecking 280GPUs, LlamaVision-20B, $30\%\downarrow$
Frequent GPU mem. management 1344GPUs, Llama-176B, $19\%\downarrow$
Void percentage
Dataloader 512GPUs, Llama-80B, $41\%\downarrow$
7.2 Case-1: Towards Stall-free Kernel Issuing
Kernel-issue stalls are among the most frequent causes of slowdowns encountered in a training cluster not dedicated exclusively to a single pre-training task. While Python runtime GC is well-known and now carefully managed by the parallel backbone in most cases, most encountered kernel-issue stalls arise from code introduced by algorithm teams to enhance LLM performance in downstream tasks.
A typical case encountered by XPUTimer is a training job of Llama20B running on 256 H800 GPUs. The developer from the algorithm team mistakenly enables the timer provided by Megatron for performance profiling of several key code segments. This profiling incurs kernel-issue stalls because it requires GPU synchronizations to obtain accurate timestamps.
Although no significant slowdown was observed in training throughput, XPUTimer successfully detects the abnormal issue latency distribution. After removing these unnecessary synchronizations by disabling the timer, the MFU of the training job improves from $41.4\%$ to $42.5\%$, representing a $2.66\%$ increase. While this slowdown is much smaller than that caused by Python runtime GC (typically exceeding $10\%$), XPUTimer could still uncover it, ensuring the training job’s performance.
In addition to the above two cases, XPUTimer also detects other kernel-issue stalls, such as unnecessary package version checking, frequent CUDA memory management within the PyTorch runtime, and others. These cases are listed in Table 3.
7.3 Case-2: Migration between Backbones
Different parallel backbones are suited to varying hardware conditions. While FSDP typically delivers ease of use and good efficiency on fewer than 1000 GPUs or relatively short sequences for LLMs (e.g., 4k), Megatron demonstrates superior scalability as the scale increases beyond 1000 GPUs or when handling extremely large sequences (e.g., 64k). In this context, algorithm teams may migrate an LLM between backbones to meet their specific demands. However, this migration process can potentially introduce anomalies.
A typical scenario encountered by XPUTimer is an anomalous MFU decline when migrating a Llama-like 80B model from FSDP (1888 H800 GPUs) to Megatron (1586 H800 GPUs with a data-parallel degree of 58, pipeline-parallel degree of 8, and tensor-parallel degree of 4). This anomaly specifically stems from a matrix layout change. The weight dimension of the LLM’s FFN layer, initially configured as $[8192\times 33936]$ during training on FSDP, changes to $[8192\times 8484]$ after migration to Megatron with a tensor parallelism degree of 4.
After migration to Megatron, this operator exhibits significantly lower FLOPS due to smaller batch size and the unfavorable 8484 layout for Tensor Cores, which require alignment to 128 bytes. In contrast, the dimension 33936 and larger batch size on FSDP meet this alignment requirement. Following XPUTimer’s diagnosis, our infrastructure team customizes a kernel that pads 8484 to 8512. Figure 13 further illustrates the FLOPS of the same operator before migration, after migration, and post-optimization guided by XPUTimer. As shown, the operator experiences a $65.3\%$ decline in FLOPS after migration, and XPUTimer successfully facilitates the performance diagnosis. From the perspective of the training job, the overall MFU increases from $27\%$ to $36\%$, reflecting a $33.3\%$ improvement.
Figure 13: The change in computation TFLOPS when migrating Llama-80B from FSDP to Megatron.
7.4 Case-3: New Algorithms and Data
Algorithm teams continually strive to enhance model performance by modifying the LLM architecture and incorporating new training data. However, this process often introduces anomalies.
Firstly, algorithm teams generally modify position embeddings (PE), activation functions (ACT), and normalization operators (NORM), while preserving the core structure of the transformer. Table 4 illustrates the detected changes in $V_{minority}$ caused by these operator modifications during daily deployments. In this table, the parallel backbone is Megatron. The “Healthy” column represents a fully optimized training job, whereas the “-PE” column reflects changes in modifying the position embeddings. Similarly, the other columns correspond to modifications of the respective operators.
Since these modified operators are less critical and not instrumented by XPUTimer, $V_{minority}$ increases proportionally with their computational complexity. Our infrastructure team leverages XPUTimer’s detection of high $V_{minority}$ to develop targeted kernel implementations. Once optimized through techniques like kernel fusion, the job’s $V_{minority}$ returns to a normal level. In this process, XPUTimer eliminates the need for manual identification, thereby accelerating anomaly diagnosis.
Table 4: Changes in detected $V_{minority}$ and corresponding TFLOPS when different minority kernels are not optimized.
Healthy -PE -PE-ACT -PE-ACT-NORM
$\mathbf{V_{minority}}$ $9\%$ $14\%$ $15\%$ $28\%$
N. TFLOPS 1 0.95 0.93 0.83
Secondly, newly filtered data is continuously incorporated into the LLM for training. XPUTimer has successfully diagnosed anomalies arising from variance in the training data. In one specific case, the algorithm team attempts to train a Llama-80B model with data containing a sequence length of 64k, while the original training script is designed for sequence lengths of 4k. XPUTimer identifies a significant anomalous decline in MFU ($41\%$) on 512 H800 GPUs, accompanied by an increase in $V_{inter}$.
After routing this anomaly to the infrastructure team, the root cause was identified in the dataloader, specifically in the attention mask generation process. When the sequence length is short, the latency incurred by mask generation is minimal. However, the complexity of mask generation scales as $O(L^{2})$, where $L$ represents the sequence length. As a result, the dataloader experiences extremely poor performance when the sequence length increases to 64k.
8 Lessons Learned and Future Work
In this section, we share the practical lessons learned during the development and deployment of XPUTimer, along with directions for future work.
8.1 Practical Usages
Algorithm teams.
Algorithm teams are dedicated to continuously integrating new innovations into the existing LLM training pipeline. They employ XPUTimer to verify that submitted training jobs meet the expected training throughput. By leveraging XPUTimer, algorithm teams can identify and resolve slowdown anomalies caused by elementary inefficient code, without requiring intervention from the infrastructure team.
Infrastructure team.
The infrastructure team is committed to optimizing GPU kernel libraries, training frameworks, and parallel backbones. With XPUTimer’s lightweight logging system, the infrastructure team can gather sufficient runtime data to analyze submitted jobs and identify new optimization opportunities. Anomalies caused by codebase modifications that cannot be resolved by the algorithm team are ultimately routed to the infrastructure team.
Operations team.
The operations team is responsible for maintaining the stability of all low-level resources. When a training job is terminated due to low-level issues, the operations team queries intra-kernel tracing information provided by XPUTimer. This allows them to identify the faulty machine without requiring a fully comprehensive test.
8.2 Holistic Diagnostics
The training stack for advancing AIGC is inherently complex. Previously, in the absence of a holistic diagnostic framework like XPUTimer, significant effort was wasted on communication among algorithm, infrastructure, and operations teams. Algorithm teams often required collaboration with the infrastructure team to address slowdowns caused by minor but inefficient codebase modifications. When low-level issues arose, the operations team had to conduct extensive low-level benchmarking tests to identify the problem, resulting in delays to the algorithm teams’ training jobs. The holistic diagnostics provided by XPUTimer streamline this process by pinpointing elementary inefficiencies in the codebase for algorithm teams, only routing complex anomalies with aggregated metrics to the infrastructure team, and supplying the operations team with valuable runtime low-level data.
8.3 Using Historical Data
Historical traces are essential for enhancing the effectiveness of anomaly detection. While runtime data is crucial, it is insufficient on its own for accurately identifying anomalies. XPUTimer detects various issues by comparing real-time data against historical data. For instance, when using issue latency distributions to detect kernel-issue stalls, XPUTimer relies on historical data from specific backbones operating on specific hardware. By collecting historical data from healthy jobs, XPUTimer can efficiently identify anomalies in newly submitted LLM training jobs. In the future, as deployment data grows, we aim to release relevant datasets to further streamline the diagnosis of LLM training.
8.4 Hardware Extensibility
Currently, XPUTimer does support the extensibility of additional hardware, particularly NPUs dedicated to DL training. Its tracing mechanism is not only backbone-agnostic but also designed for seamless extension to other NPUs. Since XPUTimer directly instruments key code segments at the Python and C++ runtime levels, extending it is straightforward, provided the relevant computation kernel is supplied by the NPU vendor.
8.5 Diagnosing Communication Slowdown
Diagnosing communication slowdowns caused by issues such as network jitter is challenging, as runtime tracing can introduce significant overhead or lack accuracy. Currently, XPUTimer utilizes a method similar to MegaScale [2], leveraging NCCL tests to identify faulty machines or switches. We are in the process of developing an eBPF-based [49] tracing tool to accurately monitor bandwidth across RDMA NICs with minimal overhead. In the future, XPUTimer will be further enhanced with these fine-grained tracing capabilities.
9 Related Work
Anomaly Diagnosis in Distributed Training.
Anomaly diagnosis in large-scale distributed deep learning training tasks has consistently been a hot research focus[2, 14, 3, 50, 51, 52]. Megascale [2] only focuses on LLM training tasks based on Megatron-LM. It identifies network-related hardware and software issues in the training process by conducting intra-host network tests and NCCL tests. Falcon [14] detects prolonged iterations using the Bayesian Online Change-Point Detection algorithm. C4D [3] modifies the Collective Communication Library to collect message statistics, such as sizes and durations of transfers, to identify the performance bottlenecks. However, these approaches primarily address communication-related performance issues and fail to encompass the anomalies across the LLM training stack. Additionally, they depend heavily on a single parallel backbone, Megatron, which limits their generality. Meanwhile, many researches[50, 51, 52] focus on task recovery. These approaches are orthogonal to XPUTimer, which could seamlessly integrate with XPUTimer.
Anomaly Diagnosis in Large-scale Datacenter.
Many research efforts[53, 54, 55, 56, 57] focus on anomaly diagnosis at different levels of the datacenter, including runtime, network, and storage. SCALENE[53] introduces an algorithm to assist Python programmers in optimizing their code by distinguishing between inefficient Python execution and efficient native execution. AND[54] is a unified application-network diagnosing system that leverages a single metric, TCP retransmissions (TCP retx), to identify network anomalies in cloud-native scenarios. AAsclepius[55] proposes a PathDebugging technique to trace fault linkages between the middle network and autonomous systems. Researchers [56] from Alibaba analyze four key factors that impact SSD failure correlations: drive models, lithography, age, and capacity. As these works address anomaly problems in specific scenarios, they are unable to resolve the challenges targeted by XPUTimer in large-scale distributed LLM training.
10 Conclusion
In this paper, we introduce XPUTimer, a real-time diagnostic framework for LLM training. By addressing critical challenges such as lightweight long-term monitoring, root cause detection, and backbone extensibility, XPUTimer provides a comprehensive solution for anomaly diagnostics across large-scale GPU clusters. With its novel intra-kernel tracing mechanism and holistic aggregated metrics, XPUTimer not only reduces diagnostic complexity but also improves the detection and resolution of non-obvious slowdowns and errors. Its deployment in real-world training clusters, spanning over 6,000 GPUs, demonstrates its efficacy in diagnosing distributed LLM training. Furthermore, XPUTimer is open-sourced to encourage broader adoption.

### v2

@@
-XPUTimer: Anomaly Diagnostics for Divergent LLM Training in GPU Clusters of Thousand-Plus Scale
+Flare: Anomaly Diagnostics for Divergent LLM Training in GPU Clusters of Thousand-Plus Scale
 Abstract
 The rapid proliferation of large language models has driven the need for efficient GPU training clusters.
-However, ensuring high-performance training in these clusters is challenging due to the complexity of software-hardware interactions and the frequent occurrence of training anomalies.
+However, it is challenging due to the frequent occurrence of training anomalies.
 Since existing diagnostic tools are narrowly tailored to specific issues, there are gaps in their ability to address anomalies spanning the entire training stack.
-In response, we introduce XPUTimer, a real-time diagnostic framework designed for distributed LLM training at scale.
-XPUTimer first integrates a lightweight tracing daemon to monitor key code segments with minimal overhead.
-Additionally, it features a diagnostic engine that employs novel intra-kernel tracing and holistic aggregated metrics to efficiently identify and resolve anomalies.
-Deployment of XPUTimer across 6,000 GPUs over eight months demonstrated significant improvements across the training stack, validating its effectiveness in real-world scenarios.
+In response, we introduce Flare, a diagnostic framework designed for distributed LLM training at scale.
+Flare first integrates a lightweight tracing daemon for full-stack and backend-extensible tracing.
+Additionally, it features a diagnostic engine that automatically diagnoses anomalies, with a focus on performance regressions.
+The deployment of Flare across 6,000 GPUs has demonstrated significant improvements in pinpointing deficiencies in real-world scenarios, with continuous operation for over eight months.
 1 Introduction
 The advent of large language models (LLMs) has revolutionized the deep learning training community, driving substantial advancements in artificial intelligence-generated content (AIGC).
-Recognizing their transformative potential to enhance user experiences, leading corporations are proactively leveraging LLMs to enhance a wide array of user-oriented services [1, 2, 3].
+Recognizing their transformative potential to enhance user experiences, leading corporations are proactively leveraging LLMs to enhance a wide array of user-oriented services [antgroup, jiangMegaScaleScaling, dongEnhancingLargescale].
 To meet the computational demands of LLM training, they construct large-scale training clusters comprising the latest GPUs interconnected via high-bandwidth links.
 Figure 1 depicts the general training stack of the large-scale training cluster in modern corporations.
-As shown, the operations team manages low-level resources, and the infrastructure team delivers training optimizations[4, 5, 6, 7], with particular emphasis on parallel backbones[8, 9, 10].
-Supported by these two teams, multiple algorithm teams focus on adapting LLMs for user-facing applications through various training methods [11, 12].
-Notably, since LLMs do not solve all problems, the training cluster also supports other deep learning jobs, such as recommendation models and their specific parallel backbone, TorchRec [13].
-Figure 1: The summarized training stack of large-scale training cluster in Ant Group, highlighting XPUTimer’s position.
-Efficient distributed LLM training of large scale requires delicate collaboration among the teams across the stack.
-However, due to the involvement of thousands of software and hardware components, training anomalies frequently arise from various layers.
-These anomalies include obvious job failures and non-obvious training slowdowns, which should be resolved by different teams.
-Specifically, the algorithm teams may unintentionally use incorrect training configurations, leading to training slowdowns.
-Meanwhile, the slowdowns may also come from unoptimized operators, which fall under the responsibility of the infrastructure team.
-Moreover, a significant number of job failures are caused by hardware errors.
-However, the current training stack lacks a dedicated diagnostic framework to address these anomalies.
-In its absence, resolving such issues requires cross-team investigations, resulting in high communication overhead and low resolution efficiency.
-In order to design a deployable diagnostic framework in the training cluster, we identify three main challenges based on real-world operational experiences.
-C-1: Designing a long-running and lightweight diagnostic framework is challenging.
-Addressing non-obvious training slowdowns and low-level communication errors requires long-term and in-depth monitoring.
-However, exhaustive tracing mechanisms, such as PyTorch’s built-in profiler, impose a high runtime memory overhead, making it unsuitable for meeting long-running requirements.
-C-2: Detecting and diagnosing the underlying root cause is challenging.
-Errors or slowdowns in LLM training often manifest with similar symptoms, such as process hangs or decreased training speed, which obscure the actual problematic machine or code.
-C-3: Providing a backbone-extensible diagnostic mechanism is challenging.
-As shown in Figure 1, at least three backbones are commonly used across the LLM community, not to mention those supporting other deep learning models.
-Making diagnostic mechanisms extensible to multiple backbones is a significant challenge.
-Faced with these challenges, previous efforts[14, 2, 3] targeting distributed LLM training fail to provide comprehensive solutions.
-This is because these approaches are narrowly designed for specific problems or scenarios, making them inadequate for addressing issues across the software-hardware stack.
-For example, FALCON [14] and C4D [3] are capable of diagnosing low-level network issues.
-However, our observations suggest that many slowdowns are caused by upper-level teams, as shown in Table 1.
-Likewise, tools such as MegaScale [2] are optimized for pre-training scenarios where teams collaborate closely and use only the Megatron backbone, limiting their adaptability to the advancing training stack.
-To this end, we present XPUTimer, a real-time and holistic anomaly diagnostic framework designed for efficient distributed LLM training in GPU clusters of thousand-plus scale.
-XPUTimer consists of two components: a per-training-process tracing daemon and a diagnostic engine.
-To collect runtime data with low overhead (C-1), the tracing daemon instruments only key code segments rather than blindly tracing all runtime data.
-Furthermore, as the code segments are carefully selected and intercepted at the level of both Python and C++ runtime, the instrumentation provides sufficient critical information for diagnosis (C-2) while remaining backbone-agnostic (C-3).
-Notably, XPUTimer is also hardware-extensible, which could provide seamless support for other NPUs.
-With the real-time data collected by the tracing daemon, the diagnostic engine detects and diagnoses anomalies, including errors and slowdowns(C-2).
-For error diagnostics, XPUTimer introduces a novel intra-kernel tracing mechanism, providing fine-grained diagnostics specifically targeting communication-related hang errors.
-Compared to a binary search using NCCL tests [5, 2], it reduces the complexity of faulty machine diagnostics from $O(\log N)$ to $O(1)$.
-For slowdown diagnostics, XPUTimer proposes holistic aggregated metrics that encompass not only commonly used macro metrics like training throughput but also novel micro metrics, such as issue latency distribution.
-With these metrics, XPUTimer can diagnose non-obvious slowdowns (e.g., $2.66\%$) in real-world workloads.
-[…]
-XPUTimer has been open-sourced at https://github.com/intelligent-machine-learning/dlrover/tree/master/xpu_timer.
-XPUTimer serves as a core component of DLRover [15], an automated distributed DL system deployed within Ant Group and supported by the LF AI & Data Foundation [16].
+As shown, the operations team manages low-level resources, and the infrastructure team delivers training optimizations[cublas2024, nccl2024, daoFlashAttentionFast, paszkePyTorchImperative], with particular emphasis on parallel backends[zhaoPyTorchFSDP, shoeybiMegatronLMTraining, rajbhandariZeROMemory].
+Supported by these two teams, various algorithm teams focus on adapting LLMs for user-facing applications through various training methods [shengHybridFlowFlexible, hanParameterEfficientFineTuning].
+Notably, the training cluster also supports other deep learning jobs, such as recommendation models and their specific parallel backend, TorchRec [ivchenkoTorchRecPyTorch].
+Figure 1: The training stack of large-scale training cluster.
+Efficient large-scale distributed LLM training requires effective coordination of software and hardware across the stack.
+However, training anomalies frequently arise from various layers.
+These anomalies include obvious job failures and non-obvious training slowdowns, which can be categorized into three types: performance regressions, job failures, and fail-slows.
+Performance regressions may result from incorrect training settings (addressed by algorithm teams) or unoptimized operators (addressed by infrastructure teams).
+In contrast, most job failures and fail-slows arise from hardware issues, which are handled by the operations team.
+One key insight in this paper is that fail-slows are sudden and easy-to-detect slowdowns due to transient issues in components, whereas performance regressions are persistent and hard-to-detect slowdowns caused by code updates or configuration drift.
+Drawing on our daily operational experiences, teams have spent significant time in cross-team collaborations to address these anomalies, often repeating efforts on similar cases.
+For instance, performance regressions caused by unnecessary synchronizations introduced by algorithm teams often require collaboration with the infrastructure team.
+Yet, eliminating these synchronizations usually only involves deleting a few lines of code in the training scripts.
+Therefore, a dedicated diagnostic framework is urgently needed to mitigate this.
+The framework first should collect sufficient runtime data for anomaly detection and potential causes identification, enabling teams to resolve them with minimal collaboration.
+To this end, runtime data must be continuously gathered from all layers of the training stack, especially from the software in the algorithm and infrastructure teams.
+Secondly, root causes should be narrowed as much as possible; otherwise, even if recurrent anomalies are routed to the right teams, they may still be unable to resolve them independently.
+To design such a diagnostic framework, we identify two main challenges.
+C-1: Designing a full-stack tracing mechanism with backend extensibility is challenging.
+The tracing mechanism collects runtime data for anomaly diagnosis.
+While full-stack tracing is needed to capture sufficient runtime data, diverse parallel backends calls for backend-extensibility, However, these requirements often conflict, as full-stack tracing usually intrudes into parallel backend codebases, reducing backend-extensibility.
+C-2: Detecting and diagnosing root causes is challenging.
+Errors or fail-slows in LLM training often show similar symptoms, such as process hangs or sudden drops in training speed.
+Worse, regressions cannot be directly confirmed by comparing macro metrics, such as current training throughput, with historical runs, since throughput declines may be expected.
+These factors make it difficult to detect anomalies and narrow down their root causes.
+Faced with these challenges, previous efforts[wuGREYHOUNDHunting, jiangMegaScaleScaling, dongEnhancingLargescale] targeting distributed LLM training fail to provide comprehensive solutions.
+This is because they are narrowly designed for specific problems or scenarios, making them inadequate for addressing issues across the software-hardware stack.
+For example, Greyhound [wuGREYHOUNDHunting], C4D [dongEnhancingLargescale], and Holmes [yaoHolmesLocalizing] can only trace and diagnose errors or fail-slows under the responsibility of operations team.
+However, they could not solve the performance regressions originated from upper-level teams.
+Besides, tools such as MegaScale [jiangMegaScaleScaling] are optimized for pre-training scenarios involving a single algorithm team and a single backend.
+It intrudes into the backend codebase to achieve full-stack tracing, but this tight coupling makes it difficult to plug into other parallel backends.
+Moreover, MegaScale only provides visualization of distributed training and relies on cross-team collaboration and manual effort to pinpoint root causes for detected anomalies.
+To this end, we present Flare, an anomaly diagnostic framework for divergent LLM training in GPU clusters at thousand-plus scale.
+It places particular emphasis on diagnosing performance regression anomalies from algorithm and infrastructure teams.
+Flare consists of two components: a per-training-process tracing daemon and a diagnostic engine.
+To collect sufficient runtime data with backend-extensibility (C-1), the tracing daemon selectively instruments key code segments in a plug-and-play manner across both Python and C++ runtimes.
+Most importantly, it leverages CPython tracing mechanisms to trace Python code segments without intruding into backbone codebases.
+It also provides an easy-to-play interface—simply adding shell environments—for extending to new backends and allowing all teams to selectively trace Python code segments.
+With real-time data collected by the tracing daemon, the diagnostic engine detects and diagnoses anomalies while narrowing their root causes (C-2).
+As for error diagnostics, Flare introduces a novel intra-kernel inspecting mechanism, providing fine-grained diagnostics specifically targeting communication-related hang errors.
+It enables $\mathcal{O}(1)$-complexity faulty machine diagnostics without exhaustive or blind searches based on NCCL tests.
+As for slowdown diagnostics, Flare proposes holistic aggregated metrics that encompass not only commonly used macro metrics like training throughput but also novel micro metrics, such as issue latency distribution.
+Importantly, the new proposed micro metrics enable Flare to detect and diagnose regressions.
+Overall, with these metrics, Flare enables teams to independently address those routed slowdowns, including subtle ones such as a $2.66\%$ regressions in real-world workloads.
+We extensively evaluated Flare in terms of its runtime overhead.
+Flare incurs an average latency overhead of only $0.43\%$ across various LLMs and backends on 1024 H800 GPUs.
+Meantime, Flare only generates just $1.5\text{MB}$ of tracing logs per GPU in a real-world model trained on 1536 H800 GPUs.
+Flare has been deployed in our training cluster at Ant Group[antgroup], utilizing over 6,000 GPUs to date.
+The greatest benefit of Flare’s deployment is that it eases the handling of recurrent anomalies.
+Flare has been open-sourced at DLRover.
+It serves as a core component of DLRover [dlrover], an automated distributed DL system deployed within Ant Group and supported by the LF AI & Data Foundation [lfaidata].
 Our contributions are as follows.
   • We highlight the urgent need for a real-time, holistic diagnostic framework capable of identifying LLM training anomalies across the entire training stack.
-  • We introduce XPUTimer, a diagnostic framework specifically designed to tackle the critical challenges of long-term monitoring, root cause diagnosis, and backbone extensibility in LLM training diagnostics.
-  • We deploy XPUTimer across more than 6,000 GPUs over an 8-month period, deriving typical case studies and practical insights from its daily operations.
-Table 1: A comprehensive analysis of common anomalies encountered in Ant Group, annotated with XPUTimer’s primary target.
+  • We introduce Flare, a diagnostic framework specifically designed to tackle the critical challenges of full-stack tracing, root cause diagnosis, and backend extensibility in LLM training diagnostics.
+  • We deploy Flare across more than 6,000 GPUs over an 8-month period, deriving typical case studies and practical insights from its daily operations.
+Table 1: A comprehensive analysis of anomalies encountered in Ant Group, with Flare’s primary target highlighted in yellow.
 Anomalies
@@
 Taxonomy
-Algorithm
-bugs
-Infrastructure
-bugs
 OS
@@
 errors
-Network
-errors
-New
-algorithms
-Unnecessary
-synchronization
-Un-optimized
-kernels
-Memory
-management
+[…]
 GPU
@@
 Symptom
-Startup crash or
-hang error
 Runtime hang
 or crash error
-Slowdown compared to historical
-jobs and prior training steps
-Slowdown compared to
+[…]
+Fail-slows compared to
 prior training steps
-Team Algorithm Intrastructure Operations Algorithm Intrastructure Operations
-Attribution Obvious Obscure Obvious
-Comment No need to diagnose XPUTimer’s main target
+[…]
 2 Background and Motivation
-In this section, we introduce the training of LLMs within corporations like Ant Group, and motivate the design of XPUTimer for anomaly diagnostics of LLM training.
 2.1 Large-Scale LLM Training Stack
-The pursuit of more powerful large language models [17, 18, 19, 20, 21, 22, 23, 24, 25, 26] has sparked an intense competition between leading corporations.
-Referring to the software-hardware stack for training LLMs in Figure 1, we delve into how LLMs are reshaping corresponding teams.
+Referring to the software-hardware stack in Figure 1, we delve into how LLMs are reshaping teams in leading corporations.
 Advancing of LLM applications.
-With the advancing of LLM algorithms [18], LLMs excel not just in natural language understanding, but also in tackling advanced tasks such as multimodal tasks  [27, 28], reasoning tasks [29, 30].
+With the advancing of LLM algorithms [deepseekmoe], LLMs excel not just in natural language understanding, but also in tackling advanced tasks such as multimodal tasks  [vit, sora], reasoning tasks [openaio1, rstar].
 For instance, customer service teams fine-tune LLMs to develop chatbots that generate accurate responses to complicated questions.
-Similarly, product development teams leverage multimodal LLMs to generate comprehensive product descriptions by integrating diverse inputs, such as product images and textual data.
-Consequently, different algorithm teams continue to innovate LLM models [23, 24, 25, 26, 18, 31, 32, 33, 34, 35, 11, 36] tailored for diverse application scenarios.
+Similarly, product development teams leverage multimodal LLMs to generate comprehensive product descriptions by integrating inputs, such as product images and textual data.
+Consequently, different algorithm teams continue to innovate LLM models tailored for diverse application scenarios.
 Advancing of training cluster.
-Training these LLMs is computationally intensive, requiring large-scale GPU clusters of thousand-plus scale [26].
-Therefore, leading corporations are continuously investing in large training clusters powered by state-of-the-art hardware [37, 38].
-These clusters not only expand in scale but also incorporate the latest GPUs [39, 40] with higher computational performance and advanced interconnect [41, 42].
+Training these LLMs is computationally intensive, requiring large-scale GPU clusters of thousand-plus scale [llama3-1].
+Therefore, leading corporations are continuously investing in large training clusters powered by state-of-the-art hardware [xai_colossus_training_cluster, meta_ai_training_infrastructure].
+These clusters not only expand in scale but also incorporate the latest GPUs [nvidia_ampere_architecture, nvidia_hopper_architecture] with higher computational performance and advanced interconnect [nvidia_nvswitch, nvidia_nvl72].
 Efficient operation of these large-scale clusters is critical to the corporations, which requires the operations team to ensure uninterrupted training performance.
-Its responsibilities include job scheduling, driver updates, hardware maintenance, etc.
 Advancing of training infrastructure.
@@
 This stack bridges the gap between algorithm teams and operations team.
-Specifically, the infrastructure team focuses on optimizing the software stack by integrating advanced operator libraries [4, 43, 5], training framework [7], and state-of-the-art model parallel backbones [8, 9, 44, 13].
-By leveraging this highly optimized training infrastructure, we unlock the full potential of LLMs, enabling innovation and scalability across our various services.
+Specifically, the infrastructure team focuses on optimizing the software stack by integrating advanced operator libraries [cublas2024, cutlass2024, nccl2024], training framework [paszkePyTorchImperative], and state-of-the-art model parallel backends [zhaoPyTorchFSDP, shoeybiMegatronLMTraining, liangTorchTitanOnestop, ivchenkoTorchRecPyTorch].
+Parallel backends are fundamental for efficiently training LLMs, enabling tensor, data, and pipeline parallelism across thousands of GPUs.
 2.2 Anomalies of Large-scale LLM Training
@@
 These issues affect both the training speed of individual training jobs and the overall utilization of training clusters, collectively termed as anomalies.
-Table 1 presents a distilled analysis of the common anomalies in our real-world cluster, broadly categorized into two primary types: errors and slowdowns.
-These anomalies often span responsibilities across multiple teams  [2, 3, 14], making them notoriously difficult to diagnose.
-Thus, resolving these anomalies demands a diagnostic framework that can effectively identify their attributions.
-However, designing such a framework is far from trivial, posing three key challenges.
-Figure 2: Slowdown of a 1024-GPU training job with Llama2-70B under various single-GPU underclocking configurations. Results are evaluated using both Megatron [9] and FSDP [8].
-The pain of triggering condition.
-Anomalies in large-scale LLM training can arise at any stage and time, as shown in the fourth row of Table 1.
-First, low-level OS or hardware issues, often attributed to the operations team, may occur randomly, causing runtime hangs or slowdowns.
-For example, Figure 2 illustrates a significant MFU decline resulting from single-GPU underclocking.
-Second, training slowdowns caused by codebase modifications may also occur randomly.
-Specifically, codebase updates on the LLM algorithm typically cause slowdowns at the start, while modifications related to the training infrastructure can result in mid-training slowdowns.
-A notable example is the slowdown caused by on-demand garbage collection (GC) [2] of Python runtime, as detailed in §5.2.2.
-In such cases, a long-term background diagnostic mechanism is required to identify these anomalies.
-However, given the variety of anomalies across the stack, it is challenging to design a lightweight diagnostic framework capable of efficiently collecting key metrics.
-The pain of obscured attribution.
-[…]
-The pain of backbone extensibility.
-With decades of advancements in model parallelism for large-scale training, a variety of divergent backbones have been developed to parallelize LLMs.
-Most LLMs are pre-trained using Megatron [9], a specialized parallel backbone designed for transformer-based LLMs.
-However, researchers continue to investigate alternative backbones [8, 10, 44] for fine-tuning pre-trained LLMs on downstream tasks or for training more complex large models, such as multimodal models.
-For instance, fully sharded data parallelism (FSDP)[8] and DeepSpeed[10] are widely utilized within our cluster for training multimodal LLMs.
-These alternatives are often selected to better align with specific training requirements.
-However, integrating the diagnostic framework into divergent parallel backbones remains challenging, as their runtime environments and programming interfaces vary significantly.
-Existing works have made efforts to build a more robust training stack for large-scale LLM training [2, 3].
-They generally operate under several assumptions: (1) training tasks rely on limited backbones like Megatron, (2) all teams work in close collaboration to pre-train a single LLM, or (3) The focus is on specific anomalies, such as communication-related issues.
-These assumptions fail to reflect reality and these works cannot solve the problems across the full stack.
-To this end, we propose XPUTimer, a cluster-wide solution deployed in Ant Group’s training cluster.
-It monitors the real-time progress of LLM training jobs and diagnoses various anomalies precisely.
-3 XPUTimer Design
-Figure 3: Architecture overview of XPUTimer.
-In this paper, we propose XPUTimer, a holistic and real-time diagnostic framework that identifies anomalies in large-scale training with low overhead and pinpoints their corresponding root causes.
-XPUTimer is designed and implemented following three principles.
-  • XPUTimer should be lightweight enough. In this case, it could run as a long-term diagnostic mechanism without compromising the training performance at any scale.
-  • XPUTimer should cover enough runtime information. In this case, it could identify the attribution precisely with detailed analysis when encountering anomalies.
-  • XPUTimer should be backbone-agnostic. In this case, it is sufficiently extensible to function as a cluster-wide solution for anomaly diagnostics, regardless of the backbones employed.
-Architecture overview.
-Figure 3 illustrates the architecture of XPUTimer, which is deployed in Ant Group’s large-scale training cluster. XPUTimer is composed of two components: the tracing daemon and the diagnostic engine.
-By automatically attaching a tracing daemon to each training process in LLM training jobs, XPUTimer enables a framework-agnostic and lightweight tracing mechanism.
-Specifically, the tracing daemon instruments only key code segments at the level of Python and C++ runtime, focusing on carefully selected Python APIs and GPU kernels.
-During instrumentation, events are injected to measure the latencies of these code segments.
-This approach draws on our experience hosting various large-scale LLM training jobs, targeting critical paths that commonly impact training efficiency.
-Then, the timing data collected by the daemons is transmitted to the diagnostic engine for holistic anomaly diagnostics.
-The engine employs a fast hang-error diagnostic method and leverages novel aggregated metrics to effectively identify slowdowns.
+Table 1 presents a distilled analysis of the common anomalies in our real-world cluster, broadly categorized into two primary types: runtime errors and slowdowns.
+Based on symptom differences, slowdowns can be further divided into two types: performance regressions (persistent slowdowns) and fail-slows (sudden and acute slowdowns).
+Based on three months of traces from a cluster with 6000+ GPUs, there are 127 errors and 135 slowdowns observed across 3047 jobs.
+The slowdowns include 78 performance regressions and 57 fail-slows.
+When such anomalies occur, they are first reported by algorithm teams, and then multiple teams coordinate and spend significant effort to resolve them.
+Specifically, for runtime errors and fail-slows, the operations team intervenes to resolve these acute and easy-to-detect job failures.
+Without timely intervention, training jobs may fail or run at very low throughput, significantly reducing training efficiency.
+For performance regressions, algorithm teams are often unfamiliar with why these regressions occur with their own code changes, so collaboration with infrastructure teams is required to resolve them.
+In a training stack like Figure 1, the situation worsens: the infrastructure team supports multiple algorithm teams, each potentially using different parallel backends for LLM training.
+As recurrent regressions arise across different algorithm teams, the infrastructure team must repeatedly expend effort to address them.
+This dilemma motivates a dedicated diagnostic framework to streamline anomaly resolution, especially performance regressions.
+To achieve this, we must collect sufficient runtime data across the training stack.
+For example, overall training throughput helps detect fail-slows, while per-operator performance reveals regressions introduced by minor software changes.
+We must also seamlessly support multiple parallel backends used by different algorithm teams.
+Finally, for detected anomalies, we aim to narrow down their root causes as much as possible, enabling them to be routed to the appropriate team and resolved independently without unnecessary cross-team collaboration.
+However, designing such a framework is far from trivial, posing two key challenges.
+The difficulty of full-stack, backend-extensible tracing.
+Diagnosing performance regressions across diverse parallel backends requires full-stack and backend-extensible tracing to collect sufficient runtime data.
+These goals often conflict.
+Full-stack tracing tends to be intrusive to each backend’s codebase, whereas extensibility demands minimal changes to the backends.
+A typical example is MegaScale [jiangMegaScaleScaling], which enables low-overhead tracing for FSDP–the native PyTorch backend–by patching the FSDP codebase.
+Extending support to other backends would require similar patches.
+A tracing design for runtime data collection in training jobs of multiple algorithm teams should avoid such backend-intrusive modifications.
+The difficulty of narrowing down attribution.
+Errors and fail-slows in Table 1 often show obvious symptoms, as the job either fail or show sudden training speed degradation.
+There are various works like C4D [dongEnhancingLargescale], Greyhound [wuGREYHOUNDHunting], and Holmes [yaoHolmesLocalizing] that are proposed to diagnose them.
+However, automatic performance regressions are rarely explored in prior work.
+Attributing them is more challenging than diagnosing fail-slows.
+First, they are harder to detect.
+Regressions are persistent slowdowns caused by software changes, and monitoring training throughput alone is insufficient.
+Algorithm teams may not notice them and may assume jobs are running normally.
+Second, linking a detected regression to the corresponding software change is difficult.
+For example, whether Python garbage collection causes a regression depends on how it is invoked.
+Thus, new metrics must be designed based on operational expertise, especially from infrastructure team, to detect and diagnose regressions.
+3 Flare Design
+Figure 2: Architecture overview of Flare.
+In this paper, we propose Flare, a diagnostic framework for diverse large-scale LLM training.
+It places particular emphasis on diagnosing performance regression automatically from algorithm and infrastructure teams.
+Figure 2 illustrates the architecture of Flare, deployed in Ant Group’s large-scale training cluster.
+Flare is composed of two components: the tracing daemon and the diagnostic engine.
+By automatically attaching a tracing daemon to each training process in LLM jobs, Flare provides a lightweight tracing mechanism.
+The daemon is backend-extensible but also enables the collection of sufficient runtime data across the entire training stack.
+For each job, runtime data such as function timing events and call stacks are intercepted for monitoring and diagnostics.
+The diagnostic engine employs a fast hang-error diagnostic method and leverages novel aggregated metrics to effectively identify slowdowns, especially regressions.
 The aggregated metrics encompass both macro-level metrics, such as training throughput, and micro-level metrics, including issue latency distribution of GPU kernels.
-Once the diagnostic engine identifies errors and slowdowns, the issues are routed to attribution teams, enabling swift resolution.
+The overall diagnostic pipeline is as follows:
+①Fast hang-error diagnosis enables error detection and analysis.
+Fail-slows are detected through macro metrics and validated by micro metrics.
+Errors and fail-slows are then routed to the operations team for independent handling.
+②Regressions are detected by monitoring the proposed micro metrics and diagnosed via Python API analysis.
+Flare routes potential regressions with narrowed root causes to the algorithm or infrastructure team.
+③The algorithm, infrastructure, and operations teams collaborate only when anomalies cannot be resolved independently by the routed team.
 4 Lightweight Selective Tracing
-To collect sufficient real-time data for anomaly diagnosis in the LLM training cluster, XPUTimer’s tracing daemon offers backbone-agnostic and lightweight full-spectrum tracing.
-Its design focuses on two key aspects: determining what information to collect and establishing how to collect it efficiently.
-4.1 Key Segment Instrumentation
-[…]
-Figure 4: Instrumented key code segments in XPUTimer.
-Therefore, XPUTimer selectively instruments code segments of key APIs and kernels to collect real-time information.
+The runtime data collection overhead primarily stems from high memory usage rather than interference with computing resources, since profiling APIs like CUPTI[cupti2024] can operate in a background thread.
+For instance, profiling a Llama-70B model trained on 512 H800 GPUs using PyTorch’s built-in profiler produces a log file of $5.5\text{GB}$ (in JSON format, compressed to $451\text{MB}$) for each training step.
+This substantial memory overhead renders such arbitrary profiling methods impractical for continuously collecting real-time data.
+Therefore, Flare selectively instruments code segments of key APIs and kernels to collect real-time information.
 This design is based on an insight into LLM training on large-scale GPUs: LLM training is predominantly dominated by a limited set of deep learning operators.
 These operators mainly include matrix multiplication and cross-GPU communication operators.
-Figure 4 presents the specific code segments instrumented by XPUTimer for efficient anomaly diagnostics.
-As shown in the figure, the instrumented code segments can be broadly categorized into two groups.
-The first category involves intercepting key API calls, including those related to Python’s garbage collection (GC), PyTorch’s dataloader, and GPU synchronization.
-These APIs are carefully selected based on empirical insights into performance issues and optimization opportunities.
-These insights are detailed in §5.2, with corresponding cases discussed in §7.
-The second category focuses on intercepting critical GPU computation and communication kernels executed at the C++ runtime level.
-These kernels, primarily provided by optimized libraries[4, 6, 46, 5], account for the majority of the workload during large-scale training.
-Additionally, there are customized kernels developed by the infrastructure team.
-Notably, the above design enables XPUTimer to support backbone-agnostic and extensible tracing capabilities.
-Extending tracing capabilities for Python-related APIs is straightforward, requiring only the configuration of the specific environment variables in the training scripts, as shown below.
-{minted}[ frame=none, obeytabs=true, framesep=0mm, baselinestretch=0.8, fontsize=, xleftmargin=1.6em, breaklines, escapeinside=||, ]shell export TRACED_PYTHON_API="torch.cuda@synchronize" Meanwhile, intercepting C++ kernels necessitates explicit registration through a C++ interface.
-This requirement is feasible, as the infrastructure team takes charge of the development of both these customized operators and XPUTimer, ensuring seamless integration and functionality.
-4.2 Timing in the Background
+In the rest of this section, we introduce the way Flare achieves full-stack and backend-extensible tracing through employing two techniques: plug-and-play instrumentation and timing with stack reconstruction.
+4.1 Plug-and-Play Instrumentation
+Figure 3: Instrumented key code segments in Flare.
+Full-stack, backend-extensible tracing requires plug-and-play instrumentation that captures the required runtime data without intrusive backend changes.
+Figure 3 shows the code segments instrumented by Flare, which are widely used in backend codebases and broadly fall into two groups.
+The first intercepts key API calls, including Python’s garbage collection (GC), PyTorch’s dataloader, and GPU synchronization.
+The second targets critical GPU computation and communication kernels executed at the C++ runtime level.
+These kernels dominate the workload in large-scale training.
+Intercepting C++ functions is straightforward with backend extensibility, as mechanisms like LD_PRELOAD can hook any dynamically linked C++ functions.
+In contrast, intercepting Python APIs is more challenging when enforcing backend extensibility.
+A common method for tracing is to provide interfaces such as Python decorators to wrap APIs in backend.
+E.g., MegaScale supports FSDP by patching PyTorch, which requires intruding into each backend for full-stack support.
+To avoid such intrusiveness for backend extensibility, Flare adopts a CPython-based intercepting mechanism.
+It maintains a list of tracing-required APIs for each backend.
+When a training job starts, Flare imports these APIs and retrieves their bytecode.
+During long training runs, it intercepts them directly using CPython’s profiling API PyEval_SetProfile based on the bytecode.
+The above design allows Flare to instrument either Python or C++ functions in a plug-and-play manner.
+For easy plug-in, Flare does not modify any backend codebase before tracing.
+For easy play, Flare exposes user-friendly interfaces for tracing new functions.
+Specifically, users only need to configure environment variables before launching jobs, such as export TRACED_PYTHON_API=‘‘torch.cuda@synchronize’’.
+Intercepting C++ functions requires explicit registration through a C++ interface.
+This is practical because the infrastructure team develops both these C++ functions and Flare, ensuring seamless integration and functionality.
+4.2 Timing with Stack Reconstruction
 With intercepted Python APIs and GPU kernels, FLARE measures their elapsed latencies, as shown in Figure 4.
@@
 For synchronous Python API calls, FLARE directly records their start and end timestamps and forwards them to the timing manager.
-For GPU kernels, which execute asynchronously, FLARE injects CUDA events[47] after an interception to record execution status.
+For GPU kernels, which execute asynchronously, FLARE injects CUDA events[cudaevents2024] after an interception to record execution status.
 These events are enqueued for further processing.
 The timing manager queries the status of the queued events in the background, avoiding any disruption to the training thread.
-Additionally, during GPU kernel interception, FLARE extracts input specifications, such as memory layout, to support subsequent anomaly diagnostics in §5.2.1.
-Figure 5: Intercepting and timing the training in the background.
-As training progresses, the timing manager proactively transmits all real-time data to XPUTimer’s diagnosis engine.
-By employing key segment instrumentation and running timing tasks in the background, XPUTimer minimizes both computing resource and memory overhead, ensuring efficient data collection for real-time anomaly diagnostics.
-A detailed evaluation of XPUTimer’s real-time overhead is provided in §6.1.
+Additionally, during GPU kernel interception, FLARE extracts input specifications, such as memory layout, to support subsequent anomaly diagnostics.
+As training progresses, the timing manager proactively streams real-time data to Flare’s diagnostic engine.
+However, since plug-and-play instrumentation uses separate methods for Python APIs and C++ functions, it omits the call stack linking them, which is essential for anomaly diagnostics.
+Fortunately, we record not only their durations but also their start and end timestamps.
+Using this information, the tracing thread reconstructs the call-stack relationships before sending the data to the diagnostic engine.
+Figure 4: Timing the training in the background.
 5 Anomaly Detection and Diagnosis
-Using the real-time data collected by the tracing daemon, XPUTimer’s diagnostic engine identifies and analyzes anomalies encountered during distributed LLM training.
-In this section, we present XPUTimer’s holistic diagnostic workflow for addressing two common anomaly symptoms: errors and slowdowns.
+In this section, we present Flare’s automated diagnosis of anomalies: runtime errors, fail-slows, and performance regressions.
+Among these, Flare pays particular attention to regressions, which have not been explored in prior work.
 5.1 Fast Runtime Error Diagnosis
-As shown in the left of Table 1, errors encountered at the beginning of a training job are typically caused by bugs in the training scripts, which can often be addressed easily by the algorithm teams and infrastructure team.
-However, diagnosing errors that occur during the training progress is more challenging and critical.
-Such errors often stem from issues like operating system crashes, GPU failures, or network disruptions, which can generally be resolved by isolating the problematic machines and restarting the training job.
+Runtime errors often stem from issues like operating system crashes, GPU failures, or network disruptions, which can generally be resolved by isolating the problematic machines and restarting the training job.
 A typical symptom associated with these errors is the hanging of the training job.
@@
 When the aforementioned errors occur, they rarely affect all training processes simultaneously.
-In this context, XPUTimer focuses on rapidly diagnosing hang errors by identifying faulty machines.
-Then, XPUTimer routes this information to the operations team, enabling the training job to restart with healthy machines.
-Specifically, XPUTimer’s diagnostic engine first detects hang errors by examining the status of tracing daemons.
-The tracing daemon operates in the background of the training thread and continuously queries events recorded during job execution.
+The training process hangs if a GPU fails.
+In this context, Flare focuses on rapidly diagnosing hang errors by identifying faulty machines.
+Then, Flare routes this information to the operations team, enabling the training job to restart with healthy machines.
+Specifically, the diagnostic engine detects hang errors by examining the status of tracing daemons.
+The daemon operates in the background of the training thread and continuously queries events recorded during job execution.
 If it fails to confirm the completion of an event within a predefined timeout interval, it proactively reports a potential hang error to the diagnostic engine.
-Similarly, if a tracing daemon does not transmit any real-time data within the specified timeout interval, the diagnostic engine also interprets this as an indication of a hang error.
+If a tracing daemon transmits none real-time data within the specified timeout interval, the diagnostic engine also interprets this as an indication of a hang error.
 After hang errors are reported, they are classified as either communication or non-communication errors.
-XPUTimer diagnoses these errors in two steps: first, a coarse-grained diagnosis through call stack analysis; and second, a fine-grained diagnosis using intra-kernel tracing
+Flare diagnoses these errors in two steps: first, a coarse-grained diagnosis through call stack analysis; and second, a fine-grained diagnosis using intra-kernel inspecting.
 Diagnosis using call stack analysis.
@@
 Figure 5 illustrates an example of hang-error diagnosis via call stack analysis.
-As shown in the left of Figure 6, when the training process of rank-00 crashes or is suspended due to a non-communication error, it halts at a call stack corresponding to a non-communication function.
-In contrast, the training processes of other ranks continue executing correctly and eventually stop at a call stack associated with a communication-related function that depends on coordination with rank-00.
-In this scenario, the machine associated with rank-00 is identified as the source of the error.
-It should be noted that, although these non-communication errors may cause direct crashes, the call stack analysis could also locate the faulty machine.
-However, communication hang errors cannot be identified through call stack analysis.
-As shown in the right of Figure 6, the training processes of all ranks terminate at the same call stack corresponding to a communication function, such as allreduce or allgather.
-In this scenario, there are no distinct differences between ranks based on call stack analysis.
-Figure 6: Diagnosing hang errors via call stack analysis.
+As shown in the left of Figure 5, when the training process of rank-0 crashes or is suspended due to a non-communication error, it halts at a call stack corresponding to a non-communication function.
+In contrast, the training processes of other ranks continue executing correctly and eventually stop at a call stack associated with a communication-related function that depends on coordination with rank-0.
+In this scenario, the machine associated with rank-0 is identified as the source of the error.
+However, communication hang cannot be identified through call stack analysis.
+As shown in the right of Figure 5, the training processes of all ranks terminate at the same call stack of a communication function, such as allreduce or allgather.
 We further investigate the symptoms of communication hang errors and obtain two observations.
@@
 Secondly, more hang errors result in an endless loop within the launched communication kernels, ultimately leading to job termination after a predefined timeout.
-To identify the unhealthy machine responsible for such errors, a straightforward solution is to perform a binary search by executing communication tests across all involved GPUs.
-This approach has a complexity of $O(\log N)$ and requires hours to pinpoint the faulty machine among thousands of GPUs [14].
-Diagnosis using intra-kernel tracing.
-Faced with this problem, XPUTimer introduces a minute-level diagnostic approach using intra-kernel tracing. This intra-kernel tracing leverages CUDA-GDB, the debugging tool for CUDA programming.
-Specifically, XPUTimer’s diagnostic engine instructs the tracing daemon to attach the halted training processes with CUDA-GDB before terminating them.
+Figure 5: Diagnosing hang errors via call stack analysis.
+To identify the faulty machine responsible for such errors, a common approach is to terminate the training process and run NCCL tests across all GPUs.
+However, in large-scale training where multiple model parallelisms–such as tensor [shoeybiMegatronLMTraining], pipeline [narayananPipeDreamGeneralized], and expert parallelism [lepikhinGshardScaling]–are combined, the NCCL tests must span all configured communication groups.
+In our experience, this exhaustive and blind search can take over half an hour at the thousand-GPU scale.
+Figure 6: Diagnosing hang errors via intra-kernel inspecting.
+Figure 7: A timeline of a distributed training job annotated with aggregated metrics used for diagnosing slowdowns in Flare.
+Diagnosis using intra-kernel inspecting.
+Faced with this problem, Flare introduces a minute-level diagnostic approach using intra-kernel inspecting with $\mathcal{O}(1)$ complexity.
+Instead of stopping the hanged training processes and blindly probing with NCCL tests, Flare fully utilizes real-time data at the error point for diagnosis.
+Specifically, the intra-kernel inspecting leverages CUDA-GDB, the debugging tool for CUDA programming.
+It instructs the tracing daemon to attach to the halted training processes using CUDA-GDB.
 Once attached, the tracing daemon executes a script capable of automatically extracting detailed communication statuses to identify unhealthy machines.
@@
 The data are split into chunks and thread blocks of adjacent ranks work together to transmit the chunks step by step.
-Thus, XPUTimer could retrieve the register values corresponding to the loop steps used for data transmission between linked ranks.
+Thus, Flare retrieves the register values corresponding to the loop steps used for data transmission between linked ranks.
 Theoretically, the connection with the minimum step reveals the related GPUs experiencing errors.
-This intra-kernel tracing process is performed in parallel across all involved GPUs.
+This intra-kernel inspecting process is performed in parallel across all involved GPUs.
 As a result, its complexity is $O(1)$, enabling completion within a few minutes.
-XPUTimer then routes the diagnostic information for detected errors to the operations team, assisting with tasks such as isolating faulty machines and restarting the training job.
+Flare then routes the diagnostic information for detected errors to the operations team, assisting with tasks such as isolating faulty machines and restarting the training job.
+Notably, this mechanism does not degrade the performance of the training job.
+The tracing daemon directly inspects the SASS code to retrieve register values.
+Thus, Flare does not require recompiling the NCCL library with debugging information enabled, which would disables several compiler optimizations.
+Moreover, Flare attaches to the training processes only after the job hangs, without interfering with normal execution.
 5.2 Aggregation for Slowdown Diagnosis
-As demonstrated in §2.2, slowdowns can be attributed to changes across the entire training stack.
-Meantime, slowdowns caused by software changes introduced by the algorithm and infrastructure teams are often subtle and challenging to detect.
-Identifying these changes typically requires comparisons across historical training jobs and prior training steps.
-In contrast, hardware changes, such as GPU underclocking or network jitter, are more apparent and can be detected solely through comparisons across training steps.
-To holistically identify these anomalies, XPUTimer aggregates real-time data collected from the tracing daemon into five primary metrics, shown in Figure 8.
-These metrics are based on the consensus that a “healthy” training pipeline should exhibit a timeline saturated with GPU kernels dedicated to either computation or communication.
-Computation kernels should achieve high FLOPS, while communication kernels are expected to utilize high bandwidth.
-Any deviations from these characteristics point to idle GPU resources, signaling potential slowdowns in training jobs.
-Of the five metrics, three are commonly used in existing works [2, 14], while the other two are newly introduced by XPUTimer.
-Figure 7: An example of diagnosing communication hang errors in a ring-allreduce kernel using intra-kernel tracing.
-Figure 8: A timeline of a distributed training job annotated with aggregated metrics used for diagnosing slowdowns in XPUTimer.
-5.2.1 Diagnosing Obvious Slowdown
-1 Training throughput for detecting slowdown.
-Training throughput is the most straightforward metric for detecting slowdowns.
-XPUTimer measures training throughput by timing the rate at which input data is consumed by the training pipeline.
-This is achieved by instrumenting the dataloader API of Pytorch.
-As a macro performance metric, training throughput directly reflects slowdowns in training efficiency through comparison to historical training jobs and between training steps of the same job.
-However, XPUTimer cannot diagnose the specific factors contributing to the slowdown.
-To address this, XPUTimer relies on the following four micro metrics to further investigate the underlying causes.
-2 FLOPS for slow critical kernel.
-XPUTimer monitors the FLOPS of instrumented critical computation kernels, leveraging timing data and input layout.
-By comparing the FLOPS of identical kernels across different ranks, XPUTimer diagnoses GPUs that exhibit poor computational performance, often caused by issues like GPU underclocking.
-Machines affected by GPU underclocking are then routed to the operations team for isolation.
-Additionally, by analyzing FLOPS, XPUTimer identifies un-optimized kernels in training jobs, particularly those with large input sizes but low FLOPS.
-These anomalies are detected without interrupting training jobs and are subsequently routed to the infrastructure team for further investigation.
-Notably, when analyzing FLOPS data, XPUTimer accounts for the impact of communication kernels that overlap with computation kernels.This ensures that computation kernels with falsely low FLOPS are not mistakenly flagged.
-3 Bandwidth for slow connection.
-XPUTimer monitors the bandwidth of communication kernels.
+To holistically identify fail-slow and regression slowdowns, Flare aggregates real-time data from the tracing daemon into five primary metrics, shown in Figure 7.
+These metrics reflect the consensus that a “health” training pipeline should exhibit a timeline saturated with GPU kernels dedicated to computation or communication.
+Computation kernels should deliver high FLOPS, while communication kernels should sustain high bandwidth.
+Deviations from these characteristics indicate idle GPU resources, signaling potential slowdowns in training jobs.
+In the rest of this section, we introduce the five metrics and explain how they are used to detect and narrow down the root causes of fail-slows and regressions.
+5.2.1 Detecting Fail-slows with Macro Metric
+1 Training throughput for detecting fail-slows.
+Fail-slows caused by low-level hardware changes, such as GPU underclocking or network jitter, are apparent and can be detected solely through comparisons across training steps.
+To this end, Flare measures training throughput by timing the rate at which input data is consumed by the training pipeline.
+It is achieved by instrumenting the dataloader API of Pytorch.
+5.2.2 Detecting Regression with Micro Metrics
+Regressions caused by software changes introduced by algorithm and infrastructure teams are often subtle and difficult to detect.
+Identifying them usually requires manual comparisons across historical training jobs.
+However, directly comparing training throughput between monitored and historical jobs does not reliably reveal regressions, since model size, input datasets, and other factors may differ significantly.
+To address this, Flare detects regressions using micro metrics.
+This is because when macro metrics do not show sudden changes, abnormalities in micro metrics often indicate potential persistent performance regressions.
+2 FLOPS and 3 bandwidth.
+FLOPS and bandwidth capture the performance of instrumented computation and communication kernels, and can be used to detect regressions.
+For example, computation kernels with large input sizes but much lower FLOPS than the theoretical value indicate a regression in computation performance.
+Flare monitors the FLOPS of instrumented critical computation kernels, leveraging timing data and input layout.
+It also monitors the bandwidth of communication kernels.
 A communication operator requires launching the communication kernels on all ranks.
-Since variations in kernel-issue timestamps exist across different ranks, XPUTimer calculates the communication bandwidth by utilizing the start and end timestamps of the final communication kernels issued across all participating ranks.
-The captured communication bandwidth is compared with offline profiled data.
-If low-bandwidth communication is detected, XPUTimer conducts a communication test using binary search to pinpoint machines experiencing issues such as network congestion.
-These slowdowns are then identified and routed to the operations team for resolution.
-5.2.2 Diagnosing Obscured Slowdown
-While the above three metrics ensure that both critical computation and communication GPU kernels operate at high performance, they do not cover the less critical operations, such as various CPU operations and element-wise activation GPU kernels.
-Meantime, XPUTimer’s selective key segment instrumentation also omits the monitoring of these operations.
-To diagnose their potential contributions to slowdowns, we further classify these not-instrumented operations into three categories: intra-step CPU operations, inter-step CPU operations, and minority GPU kernels.
+Since variations in kernel-issue timestamps exist across different ranks, Flare calculates the communication bandwidth by utilizing the start and end timestamps of the final communication kernels issued across all participating ranks.
+Notably, when analyzing FLOPS and bandwidth to detect regressions, Flare accounts for the overlap of communication and computation kernels, which is common in training MoE-based LLMs [deepseek-aiDeepSeekV3Technical].
+This ensures that computation kernels with falsely low FLOPS are not mistakenly flagged.
+While FLOPS and bandwidth ensure that both critical computation and communication GPU kernels operate at high performance, they do not cover the less critical operations, such as various CPU operations and element-wise activation GPU kernels.
+Meantime, Flare’s tracing also omits the monitoring of these operations.
+To detect their potential contributions to regressions, we further classify these not-instrumented operations into three categories: intra-step CPU operations, inter-step CPU operations, and minority GPU kernels.
 Intra-step CPU operations and inter-step CPU operations differ due to their occurrences within the timeline of training steps.
 Minority GPU kernels refer to those GPU kernels that often occupy little GPU computation resources.
-Specifically, two metrics are introduced for the diagnostics: issue latency distribution for intra-step CPU operations and void percentage for inter-step CPU operations and minority GPU kernels.
-Figure 9: Runtime overhead in terms of latency with various models, backbones, and number of GPUs.
+Specifically, two metrics are introduced for regression detection: issue latency distribution for intra-step CPU operations and void percentage for inter-step CPU operations and minority GPU kernels.
 4 Issue latency distribution for kernel-issue stall.
-In a well-optimized parallel backbone, only the necessary intra-step CPU operations for launching GPU kernels or coordinating the training processes are expected.
+In a well-optimized parallel backend, only the necessary intra-step CPU operations for launching GPU kernels or coordinating the training processes are expected.
 However, algorithm teams may inadvertently introduce unnecessary GPU synchronizations when modifying the LLM model.
-Meantime, certain function calls, such as GC [2, 9], may be implicitly triggered by the Python runtime.
+Meantime, certain function calls, such as GC [jiangMegaScaleScaling, shoeybiMegatronLMTraining], may be implicitly triggered by the Python runtime.
 These intra-step CPU operations can occur repeatedly during the model’s forward pass, bringing considerable overhead.
-In such cases, these operations cause an anomaly known as a kernel-issue stall, leading to GPU idle time within the training step.
+In such cases, these operations cause a regression anomaly known as a kernel-issue stall, leading to GPU idle time within the training step.
 4
@@
 As all ranks wait for the completion of communication kernels, the kernel issue of follow-up kernels is stalled and not overlapped with GPU computation.
-When such unnecessary synchronization occurs repeatedly across the model’s forward pass, it ultimately results in a slowdown of the training speed.
+When such unnecessary synchronization occurs repeatedly across the model’s forward pass, it ultimately results in a regression of the training speed.
 Originally, detecting these anomalies of kernel-issue stall requires investigating the aggregated timeline with much human effort.
-Faced with this issue, XPUTimer proposes a new metric, named issue latency distribution, for diagnosing this issue without human intervention.
+Faced with this issue, Flare proposes a new metric, named issue latency distribution, for diagnosing this issue without human intervention.
 Kernel-issue latency is defined as the time elapsed between the kernel’s issue timestamp and the start timestamp of its execution on the GPU.
-Based on our observation of anomalies of kernel-issue stall, the kernel-issue latencies of unhealthy training jobs should be much shorter than those of a healthy training job.
-By monitoring runtime issue latency distribution, and comparing it with the historical data, XPUTimer could identify the anomaly of kernel-issue stall.
-Then, XPUTimer routes them to algorithm and infrastructure teams for resolution, as they are commonly software issues.
-§6.3 and §7.2 demonstrate the effectiveness of issue latency distribution in diagnosing kernel-issue stalls, even when the slowdown is minimal.
+Based on our observation of regression due to kernel-issue stall, the kernel-issue latencies of unhealthy training jobs should be much shorter than those of a healthy training job.
+Ahead of deployment, Flare learns healthy kernel issue distributions from historical data for each backend type and cluster scale.
+Flare use the maximum Wasserstein distance [ramdas2017wasserstein] between these healthy distributions as a threshold.
+At runtime, Flare compares the collected data against the healthy distribution by computing the Wasserstein distance between them.
+A warning is triggered when the distance exceeds the learned threshold.
 5 Void percentage for other un-covered operations.
 While the tracing daemon only instruments the critical operators, inter-step CPU operations and minority GPU kernels both manifest as empty time slots in the visualized timeline, as shown in Figure 7.
-Consequently, XPUTimer introduces a metric, termed the void percentage, to identify slowdowns caused by these factors.
-As for inter-step CPU operations, as depicted by 5–2 in Figure 8, XPUTimer measures the latency between the last kernel preceding the dataloader and the first kernel following the same dataloader.
-XPUTimer then computes the void percentage for inter-step CPU operations using the following equation:
+Consequently, Flare introduces a metric, termed the void percentage, to identify slowdowns caused by these factors.
+As for inter-step CPU operations, as depicted by 5–2 in Figure 7, Flare measures the latency between the last kernel preceding the dataloader and the first kernel following the same dataloader.
+Flare then computes the void percentage for inter-step CPU operations using the following equation:
 $V_{inter}=T_{inter}\ /\ T_{step}$ (1)
 where $T_{inter}$ represents the latency associated with inter-step CPU operations, and $T_{step}$ denotes the total latency of the training step.
-As for minority GPU kernels, as shown by 5–1 in Figure 8, XPUTimer first automatically detects empty slots where GPU kernels are launched but remain un-executed.
-These empty slots signify that the GPUs are occupied by kernels outside the scope of XPUTimer’s tracing mechanism.
-XPUTimer subsequently accumulates these slots for each training step and computes the void percentage using the following equation:
+As for minority GPU kernels, as shown by 5–1 in Figure 7, Flare first automatically detects empty slots where GPU kernels are launched but remain un-executed.
+These empty slots signify that the GPUs are occupied by kernels outside the scope of Flare’s tracing mechanism.
+Flare subsequently accumulates these slots for each training step and computes the void percentage using the following equation:
 $V_{minority}=T_{minority}\ /\ (T_{step}-T_{inter})$ (2)
 where $T_{minority}$ is the latency of all minority GPU kernels.
-When the void percentages ($V_{inter}$ and $V_{minority}$) surpass the predefined thresholds for a specific parallel backbone, XPUTimer annotates the training job with potential slowdowns attributed to inter-step CPU operations or minority GPU kernels.
-Then, XPUTimer notifies the algorithm and infrastructure team for further investigation.
+When the void percentages ($V_{inter}$ and $V_{minority}$) surpass the predefined thresholds for a specific parallel backend, Flare annotates the training job with potential regressions attributed to inter-step CPU operations or minority GPU kernels.
+5.2.3 Diagnosing Root Causes for Fail-slows
+While changes in training throughput between steps of the same job indicate fail-slows, Flare cannot directly diagnose the specific factors behind these slowdowns.
+Following prior work, Flare investigates the root causes using two micro metrics mentioned above: FLOPS and bandwidth.
+By comparing the FLOPS of identical kernels across different ranks, Flare diagnoses GPUs that exhibit poor computational performance, often caused by issues like GPU underclocking.
+Machines affected by GPU underclocking are then routed to the operations team for isolation.
+The captured communication bandwidth is compared with offline profiled data.
+If low-bandwidth communication is detected, Flare conducts a communication test using binary search to pinpoint machines experiencing issues such as network congestion.
+These slowdowns are then identified and routed to the operations team for resolution.
+5.2.4 Diagnosing Root Causes for Regressions
+Pinpointing the root causes of regressions in computation and communication kernels is straightforward.
+Flare directly forwards the traced data, such as input shape and layout, to the infrastructure team for resolution.
+When Flare detects regressions related to kernel-issue stalls, it further narrows down the root causes by analyzing the invocation of relevant Python APIs recorded by the tracing daemon.
+Specifically, Flare checks for APIs such as Python GC invoked just before communication kernels with abnormal issue distributions.
+The same approach applies to regressions detected by void percentage.
+If relevant APIs are found, Flare directly pinpoints the potential root causes and first routes them to the corresponding algorithm teams.
+If the algorithm teams cannot resolve them independently, or if no relevant Python APIs are found, Flare forwards the regressions to the infrastructure team for further investigation.
+In this case, both algorithm and infrastructure teams validates and resolves regressions with the root causes already narrowed down.
 6 Evaluation
-In this section, we present experiments to demonstrate XPUTimer’s effectiveness from various perspectives.
-Specifically, we first evaluate XPUTimer’s runtime overhead in terms of latency and memory, showcasing its lightweight nature for runtime anomaly diagnosis.
-Then, we assess the effectiveness of XPUTimer’s novel diagnostic mechanisms: intra-kernel tracing for communication hang error diagnosis and issue distribution metric for kernel-issue stall diagnosis.
+Flare consists of $11.4K$ lines of code: 2,197 lines in Python, 7,447 lines in C++, and the rest in supporting code to facilitate its usage.
+Beyond the automatic diagnostic workflow described above, Flare also provides rich information to assist manual optimizations, e.g., visualized distributed training timeline.
+In this section, we present experiments to demonstrate Flare’s effectiveness from various perspectives.
+6.1 Functionality Comparison
+Since most related works [jiangMegaScaleScaling, dongEnhancingLargescale, wuGREYHOUNDHunting] are closed-source or only partially open-sourced, we begin with a functionality comparison between Flare and them, followed by in-depth evaluations of Flare.
+Table 2 summarizes the comparison.
+Firstly, in terms of tracing mechanisms, only MegaScale and Flare provide full-stack tracing.
+Other works, such as C4D, Greyhound, and Holmes, trace only communication or computation kernels.
+However, due to different design assumptions, Flare ensures backend extensibility, whereas MegaScale requires intrusive modifications to backend codebases, e.g., patching FSDP before tracing.
+Secondly, in diagnostics, only Flare offers automated detection of performance regression anomalies, mainly from algorithm and infrastructure teams, enabled by the proposed micro metrics.
+While MegaScale supports full-stack tracing, it only provides distributed visualization for manual regression investigation.
+C4D, Greyhound, and Holmes diagnose only errors and fail-slows.
+Although Flare also detects these anomalies, it does not enhance their diagnostic mechanisms beyond providing intra-kernel inspection for fast communication hang-error diagnosis.
+In a nutshell, Flare is the first diagnostic framework to combine full-stack tracing with backend extensibility and provide automated regression anomaly diagnostics.
+Table 2: Functionality comparison between Flare and other existing works. “Comm.” denotes communication.
+Category Features MegaScale C4D Greyhound Flare
+User experience Full-stack tracing
+LightCayenneBackend-extensible
+Easy-to-play interfaces
+Automated diagnostics
+with aggregated metrics
+Distributed visualization
+Hang error Non-comm. hang
+Comm. hang $\geq 30min$ $\geq 30min$ $\leq 5min$
+Slowdown Critical kernels
+Overlapping of
+Comp. and Comm.
+Comm. kernels
+Kernel-issue stall Only GC
+Less critical operations
 6.2 Runtime Overhead
-We evaluate XPUTimer on four parallel backbones: Megatron [9], FSDP [8], DeepSpeed [10], and TorchRec [13].
+Figure 8: Runtime overhead in terms of latency with various models, backends, and number of GPUs.
+We evaluate Flare on four parallel backends: Megatron [shoeybiMegatronLMTraining], FSDP [zhaoPyTorchFSDP], DeepSpeed [rajbhandariZeROMemory], and TorchRec [ivchenkoTorchRecPyTorch].
 Among these, Megatron, FSDP, and DeepSpeed are widely used for LLM training, while TorchRec is employed for training large recommendation models within Ant Group.
@@
 The latency overhead experiment is conducted on 1,024 H800 GPUs deployed across 128 servers with RoCE connectivity.
-Figure 9 presents the latency overhead introduced by XPUTimer with various models, backbones, and number of GPUs.
-As shown, XPUTimer incurs a latency overhead of $0.43\%$ for three LLM training backbones and $1.02\%$ for TorchRec.
+We directly compare Flare with the original execution, as it serves as the lower bound.
+As shown, Flare incurs a latency overhead of $0.43\%$ for three LLM training backends and $1.02\%$ for TorchRec.
+Across other GPUs like A100, Flare shows consistent results.
+We also compared the runtime overhead of Flare with existing works [jiangMegaScaleScaling, wuGREYHOUNDHunting].
+Compared with MegaScale, Flare incurs similar runtime overhead, as both selectively trace key code segments.
+The main difference is that Flare avoids intruding into backend codebases, achieving better backend extensibility.
+Originally, Greyhound only traces start timestamps of communication kernels.
+We extend Greyhound’s tracing mechanism to support full-stack tracing.
+In this case, Greyhound incurs unacceptable overhead, reaching $35\%$ latency with Llama-8B trained on just 8 GPUs.
+This is because Greyhound’s tracing mechanism is tailored for fail-slows and does not cover regressions.
 The memory overhead experiment is conducted on two setups, which are 16 A100 GPUs on 2 nodes and 1536 H800 GPUs on 192 nodes.
-We compare XPUTimer with Torch Full, Torch w/o Stack, Torch w/o Layout&Stack, and XPUTimer.
+The testbed slightly differs from that used in the latency overhead evaluation, as we conducted the experiments on available idle GPUs due to heavy load on our training cluster.
+We compare Flare with Torch Full, Torch w/o Stack, Torch w/o Layout&Stack, and Flare.
 Torch w/o Stack refers to using the PyTorch builtin profiler with stack tracing disabled, while Torch w/o Layout&Stack further disables matrix layout tracing.
-[…]
-From the above results, we can conclude that, regardless of the GPU scale, parallel backbone, parallel strategy, or model type used, XPUTimer consistently maintains an extremely low runtime overhead in terms of both latency and memory.
-The lightweight selective tracing facilitates XPUTimer’s deployment within our large training cluster, serving as a diagnostic framework for diverse training jobs.
-Figure 10: Memory consumption of dumped logs per GPU per step using the PyTorch profiler and XPUTimer while training a Llama-70B model on 16 A100 GPUs.
-6.2 Effectiveness of Intra-kernel Tracing
-We evaluate the intra-kernel tracing mechanism on 16 A100 GPUs across two servers with RoCE connectivity.
+Figure 9 shows the memory overhead results on 16 A100 GPUs.
+Flare consumes only $0.39\%$, $1.76\%$, and $2.48\%$ of memory overhead for the respective configurations of PyTorch profiler.
+Flare generates a maximum of $0.78\text{MB}$ of tracing logs per GPU.
+Besides, in a real-world Llama-20B training job on 1536 H800 GPUs, Flare generated only a $1.5\text{MB}$ tracing log per GPU.
+TorchRec is omitted from this experiment, as monitoring a recommendation model generates minimal logs.
+From the above results, Flare consistently maintains an extremely low runtime overhead in terms of both latency and memory.
+The lightweight selective tracing facilitates Flare’s deployment within our training cluster, serving as a diagnostic framework for diverse training jobs.
+Figure 9: Memory consumption of dumped logs per GPU per step using the PyTorch profiler and Flare while training a Llama-70B model on 16 A100 GPUs.
+6.3 Effectiveness of Intra-kernel Inspecting
+We evaluate the intra-kernel inspecting mechanism on 16 A100 GPUs across two servers with RoCE connectivity.
 Given that most communication kernels are ring-based, this experiment focuses on evaluating ring-allreduce.
-Specifically, we customize the training script composed solely of communication kernels, with one GPU intentionally suspended to simulate a hang error caused by communication issues.
+We customize the training script composed solely of communication kernels, with one GPU intentionally suspended to simulate a hang error caused by communication issues.
 Figure 10 illustrates the pinpointing latencies for intra- and inter-server communication.
-The figure presents the latency results for three communication protocols [48] and cross-node configurations.
-As shown, XPUTimer requires $29.4\text{\textasciitilde}309.2s$ to detect erroneous GPUs across different scenarios.
-Among the protocols, XPUTimer performs best when the SIMPLE protocol is used for communication.
-This is because, with the SIMPLE protocol, XPUTimer only needs to scan the first thread of each thread block to check the steps, whereas the other two protocols require scanning the entire thread block.
-When comparing intra-server and inter-server results, XPUTimer performs better when the ring-allreduce operation spans multiple servers.
+The figure presents the latency results for three communication protocols [NVIDIA_NCCL_User_Guide] and cross-node configurations.
+As shown, Flare requires $29.4\text{\textasciitilde}309.2s$ to detect erroneous GPUs across different scenarios.
+Among the protocols, Flare performs best when the SIMPLE protocol is used for communication.
+This is because, with the SIMPLE protocol, Flare only needs to scan the first thread of each thread block to check the steps, whereas the other two protocols require scanning the entire thread block.
+When comparing intra-server and inter-server results, Flare performs better when the ring-allreduce operation spans multiple servers.
 This is because intra-server GPUs are connected via NVLink, whereas inter-server GPUs communicate through NICs.
 Communication kernels launched over NICs involve fewer thread blocks, as NICs have fewer internal links compared to NVLink.
-As a result, XPUTimer scans fewer thread blocks for error diagnosis in inter-server scenarios.
-In summary, the intra-kernel tracing mechanism can detect erroneous GPUs in a maximum of $309.2s$.
-Notably, as the complexity of intra-kernel tracing is $O(1)$, these results remain consistent regardless of scale.
+As a result, Flare scans fewer thread blocks for error diagnosis in inter-server scenarios.
+In summary, the intra-kernel inspecting mechanism can detect erroneous GPUs in a maximum of $309.2s$. Notably, as the complexity of intra-kernel inspecting is $O(1)$, these results remain constant.
 Figure 10: Latency for pinpointing the erroneous GPUs causing a hang error in ring-allreduce with different protocols.
-Figure 12: Issue distribution across all communication kernels for a Llama-20B model trained with Megatron and 256 GPUs, including the overall CDF and the CDFs for each type of communication kernel, respectively.
+Figure 11: Issue distribution for a Llama-20B model trained with Megatron and 256 GPUs, including the overall CDF and the CDFs for each type of communication kernel, respectively.
 6.4 Effectiveness of Issue Latency Distribution
@@
 In the Unhealthy-Sync scenario, an unintended GPU synchronization call is added within the transformer block, leading to repetitive GPU synchronizations during the model’s forward pass.
-In the Healthy scenario, GC is efficiently managed by the parallel backbone, and no unnecessary synchronizations are introduced.
+In the Healthy scenario, GC is efficiently managed by the parallel backend, and no unnecessary synchronizations are introduced.
 As shown in the figure, the issue latency distribution patterns align with our claim in §5.2.2.
@@
 7 Deployment & Case Studies
-In this section, we further demonstrate XPUTimer’s effectiveness in diagnosing LLM training anomalies through statistical analysis of its cluster-wide deployment and detailed case studies of various anomalies.
 7.1 Cluster-wide Deployment
-XPUTimer has been deployed within a training cluster with 6,000 GPUs for over 8 months.
+Flare has been deployed and running continuously in a training cluster with 6,000 GPUs for over eight months.
 During this period, it is responsible for monitoring, detecting, and diagnosing training jobs for various deep learning models, especially large-scale distributed LLM training.
-Its diagnostic capabilities enable the algorithm teams, infrastructure team, and operations team to seamlessly enhance model training efficiency and improve training cluster utilization.
-Table 2: Typical errors detected by XPUTimer.
+Table 3: Typical errors detected by Flare.
 Taxonomy Details Numbers Mechanism
@@
 RoCE issue 17
-Table 2 presents a subset of error anomalies detected by XPUTimer.
-As shown in the table, XPUTimer effectively identifies OS- and hardware-related anomalies, including crashes, hangs, and slowdowns caused by such issues.
-While these issues are typically conspicuous and could be detected by existing methods based on noticeable training interruptions [9, 14], XPUTimer’s novelty lies in providing richer error information through techniques like intra-kernel tracing.
+7.2 Errors
+Table 3 presents a subset of error and fail-slow anomalies detected by Flare.
+As shown in the table, Flare effectively identifies OS- and hardware-related anomalies, including crashes, and hangs.
+While these issues are typically conspicuous and could be detected by existing methods based on noticeable training interruptions [shoeybiMegatronLMTraining, wuGREYHOUNDHunting], Flare’s novelty lies in providing richer error information through techniques like intra-kernel inspecting.
 Such runtime information helps ease and accelerate the attribution process of low-level issues for operations teams.
-Moreover, the true value of XPUTimer lies in its ability to detect anomalies caused by both algorithm teams and the infrastructure team.
-Table 3 summarizes slowdowns diagnosed by XPUTimer using its aggregated metrics.
-In the table, anomalies newly identified by XPUTimer, as compared to prior works [2], are highlighted in bold.
-The following subsections provide a detailed discussion of typical cases.
-Table 3: Slowdowns diagnosed by XPUTimer, with “Details” showing training job specifics and associated MFU decline.
+7.3 Fail-slows and Performance Regressions
+Table 4 summarizes fail-slow and regression slowdowns diagnosed by Flare using its aggregated metrics.
+We omit evaluation of Flare in fail-slow diagnostics, as it employs similar mechanisms to prior works [dongEnhancingLargescale, wuGREYHOUNDHunting].
+To evaluate regression diagnostics, we collect 113 real-world training jobs submitted within a week in a 6,000-GPU cluster.
+During this period, the cluster maintained an average GPU usage of 80%.
+The jobs included popular LLMs, multi-modal LLMs, and recommendation models.
+We first diagnose them with Flare, then validate results against human-labeled ground truth.
+It successfully diagnosed 9 true regressions based on issue latency distribution and void percentage, with only 2 false ones.
+This corresponds to a false positive rate of $1.9\%$ and a true positive diagnostic accuracy of $81.8\%$.
+We further investigated the 2 false positives.
+The first was a multi-modal LLM trained with FSDP, where input data contained images of varying resolutions.
+This caused imbalanced computation across ranks, leading to abnormal issue latency distribution.
+The second was a recommendation model using CPU-based embeddings.
+Its void percentage was higher than that of GPU-based ones, causing misclassification.
+Although these 2 cases were falsely diagnosed as regressions, they highlight opportunities to refine Flare’s micro-metrics with historical data.
+For example, relaxing latency distribution thresholds for imbalanced multi-modal inputs and adjusting void percentage thresholds for CPU-based embedding models could reduce such false positives.
+After these refinements, Flare no longer misclassifies these job types in real-world cluster deployments.
+Now, we present several typical diagnosed regressions.
+Table 4: Fail-slows and regressions diagnosed by Flare, with “Details” showing training job specifics and associated MFU decline. We show the diagnosed regression anomalies by Flare in bold.
 Metric Attribution Details
 FLOPS GPU underclocking 480 GPUs, Llama-65B, $14\%\downarrow$
-Backbone migration 1856 GPUs, Llama-80B, $33.3\%\downarrow$
+Backend migration 1856 GPUs, Llama-80B, $33.3\%\downarrow$
 Bandwidth
@@
 Dataloader 512GPUs, Llama-80B, $41\%\downarrow$
-7.2 Case-1: Towards Stall-free Kernel Issuing
+7.3.1 Case-1: Towards Stall-free Kernel Issuing
 Kernel-issue stalls are among the most frequent causes of slowdowns encountered in a training cluster not dedicated exclusively to a single pre-training task.
-While Python runtime GC is well-known and now carefully managed by the parallel backbone in most cases, most encountered kernel-issue stalls arise from code introduced by algorithm teams to enhance LLM performance in downstream tasks.
-A typical case encountered by XPUTimer is a training job of Llama20B running on 256 H800 GPUs.
+While Python runtime GC is well-known and now carefully managed by the parallel backend in most cases, most encountered kernel-issue stalls arise from code introduced by algorithm teams to enhance LLM performance in downstream tasks.
+A typical case encountered by Flare is a training job of Llama-20B running on 256 H800 GPUs.
 The developer from the algorithm team mistakenly enables the timer provided by Megatron for performance profiling of several key code segments.
 This profiling incurs kernel-issue stalls because it requires GPU synchronizations to obtain accurate timestamps.
-Although no significant slowdown was observed in training throughput, XPUTimer successfully detects the abnormal issue latency distribution.
-After removing these unnecessary synchronizations by disabling the timer, the MFU of the training job improves from $41.4\%$ to $42.5\%$, representing a $2.66\%$ increase.
-While this slowdown is much smaller than that caused by Python runtime GC (typically exceeding $10\%$), XPUTimer could still uncover it, ensuring the training job’s performance.
-In addition to the above two cases, XPUTimer also detects other kernel-issue stalls, such as unnecessary package version checking, frequent CUDA memory management within the PyTorch runtime, and others.
-These cases are listed in Table 3.
-7.3 Case-2: Migration between Backbones
-Different parallel backbones are suited to varying hardware conditions.
-While FSDP typically delivers ease of use and good efficiency on fewer than 1000 GPUs or relatively short sequences for LLMs (e.g., 4k), Megatron demonstrates superior scalability as the scale increases beyond 1000 GPUs or when handling extremely large sequences (e.g., 64k).
-In this context, algorithm teams may migrate an LLM between backbones to meet their specific demands.
-However, this migration process can potentially introduce anomalies.
-A typical scenario encountered by XPUTimer is an anomalous MFU decline when migrating a Llama-like 80B model from FSDP (1888 H800 GPUs) to Megatron (1586 H800 GPUs with a data-parallel degree of 58, pipeline-parallel degree of 8, and tensor-parallel degree of 4).
-This anomaly specifically stems from a matrix layout change.
-The weight dimension of the LLM’s FFN layer, initially configured as $[8192\times 33936]$ during training on FSDP, changes to $[8192\times 8484]$ after migration to Megatron with a tensor parallelism degree of 4.
+Although no significant regression was observed in training throughput, Flare successfully detects the abnormal issue latency distribution.
+After removing these unnecessary synchronizations by disabling the timer, the MFU of the training job persistently improves from $41.4\%$ to $42.5\%$, representing a $2.66\%$ increase.
+Such regression is minor, which could not be directly observed through macro metrics like training throughput in a large training cluster with a lot of noise.
+However, there is an obvious drift in micro metrics like issue latency distribution, enabling Flare to detect such obsecured regression anomalies.
+In addition to the above two cases, Flare also detects other kernel-issue stalls, such as unnecessary package version checking, frequent CUDA memory management within the PyTorch runtime, and others.
+7.3.2 Case-2: Migration between Backends
+Different parallel backends are suited to varying hardware conditions.
+In this context, algorithm teams may migrate an LLM between backends to meet their specific demands.
+However, this migration process can potentially introduce regressions.
+A typical scenario encountered by Flare is an anomalous MFU decline when migrating a Llama-like 80B model from FSDP (1888 H800 GPUs) to Megatron (1586 H800 GPUs with a data-parallel degree of 58, pipeline-parallel degree of 8, and tensor-parallel degree of 4).
+This regression specifically stems from a matrix layout change.
+The weight dimension of the LLM’s FFN (feed forward network) layer, initially configured as $8192\times 33936]$ during training on FSDP, changes to $8192\times 8484]$ after migration to Megatron with a tensor parallelism degree of 4.
 After migration to Megatron, this operator exhibits significantly lower FLOPS due to smaller batch size and the unfavorable 8484 layout for Tensor Cores, which require alignment to 128 bytes.
 In contrast, the dimension 33936 and larger batch size on FSDP meet this alignment requirement.
-Following XPUTimer’s diagnosis, our infrastructure team customizes a kernel that pads 8484 to 8512.
-Figure 13 further illustrates the FLOPS of the same operator before migration, after migration, and post-optimization guided by XPUTimer.
-As shown, the operator experiences a $65.3\%$ decline in FLOPS after migration, and XPUTimer successfully facilitates the performance diagnosis.
+While the algorithm team does not report this regression, Flare detected and routed it to our infrastructure team.
+Following Flare’s diagnosis, our infrastructure team customizes a kernel that pads 8484 to 8512.
+Figure 12 illustrates the FLOPS of the same operator before migration, after migration, and post-optimization guided by Flare.
+As shown, the operator experiences a $65.3\%$ decline in FLOPS after migration, and Flare successfully facilitates the performance diagnosis.
 From the perspective of the training job, the overall MFU increases from $27\%$ to $36\%$, reflecting a $33.3\%$ improvement.
 Figure 12: The change in computation TFLOPS when migrating Llama-80B from FSDP to Megatron.
-7.4 Case-3: New Algorithms and Data
-Algorithm teams continually strive to enhance model performance by modifying the LLM architecture and incorporating new training data. However, this process often introduces anomalies.
+7.3.3 Case-3: New Algorithms and Data
+Algorithm teams continually strive to enhance model performance by modifying the LLM architecture and incorporating new training data.
+However, this often introduces regressions.
 Firstly, algorithm teams generally modify position embeddings (PE), activation functions (ACT), and normalization operators (NORM), while preserving the core structure of the transformer.
 Table 5 illustrates the detected changes in $V_{minority}$ caused by these operator modifications during daily deployments.
-In this table, the parallel backbone is Megatron.
-The “Healthy” column represents a fully optimized training job, whereas the “-PE” column reflects changes in modifying the position embeddings.
+In this table, the parallel backend is Megatron.
+The Healthy column represents a fully optimized training job, whereas the -PE column reflects changes in modifying the position embeddings.
 Similarly, the other columns correspond to modifications of the respective operators.
-Since these modified operators are less critical and not instrumented by XPUTimer, $V_{minority}$ increases proportionally with their computational complexity.
-Our infrastructure team leverages XPUTimer’s detection of high $V_{minority}$ to develop targeted kernel implementations.
+Since these modified operators are less critical and not instrumented by Flare, $V_{minority}$ increases proportionally with their computational complexity.
+Our infrastructure team leverages Flare’s detection of high $V_{minority}$ to develop targeted kernel implementations.
 Once optimized through techniques like kernel fusion, the job’s $V_{minority}$ returns to a normal level.
-In this process, XPUTimer eliminates the need for manual identification, thereby accelerating anomaly diagnosis.
-Table 4: Changes in detected $V_{minority}$ and corresponding TFLOPS when different minority kernels are not optimized.
+In this process, Flare eliminates the need for manual identification, thereby accelerating anomaly diagnosis.
+Secondly, newly filtered data is continuously incorporated into the LLM for training.
+Flare has successfully diagnosed anomalies arising from variance in the training data.
+In one specific case, the algorithm team attempts to train Llama-80B with data containing a sequence length of 64k, while the original training script is for sequence lengths of 4k.
+Flare identifies a significant anomalous decline in MFU ($41\%$) on 512 H800 GPUs, accompanied by an increase in $V_{inter}$.
+After routing this anomaly with the root cause identified in the dataloader to the algorithm team, they validated that the regression was due to the attention mask generation process within the dataloader.
+When the sequence length is short, the latency incurred by mask generation is minimal.
+However, the complexity of mask generation scales as $O(L^{2})$, where $L$ represents the sequence length.
+As a result, the dataloader experiences extremely poor performance when the sequence length increases to 64k.
+Table 5: Changes in detected $V_{minority}$ and normalized TFLOPS when different minority kernels are not optimized.
 Healthy -PE -PE-ACT -PE-ACT-NORM
@@
 N. TFLOPS 1 0.95 0.93 0.83
-Secondly, newly filtered data is continuously incorporated into the LLM for training.
-XPUTimer has successfully diagnosed anomalies arising from variance in the training data.
-In one specific case, the algorithm team attempts to train a Llama-80B model with data containing a sequence length of 64k, while the original training script is designed for sequence lengths of 4k.
-XPUTimer identifies a significant anomalous decline in MFU ($41\%$) on 512 H800 GPUs, accompanied by an increase in $V_{inter}$.
-After routing this anomaly to the infrastructure team, the root cause was identified in the dataloader, specifically in the attention mask generation process.
-When the sequence length is short, the latency incurred by mask generation is minimal.
-However, the complexity of mask generation scales as $O(L^{2})$, where $L$ represents the sequence length.
-As a result, the dataloader experiences extremely poor performance when the sequence length increases to 64k.
-8 Lessons Learned and Future Work
-In this section, we share the practical lessons learned during the development and deployment of XPUTimer, along with directions for future work.
+8 Experience
 8.1 Practical Usages
-Algorithm teams.
-Algorithm teams are dedicated to continuously integrating new innovations into the existing LLM training pipeline.
-They employ XPUTimer to verify that submitted training jobs meet the expected training throughput.
-By leveraging XPUTimer, algorithm teams can identify and resolve slowdown anomalies caused by elementary inefficient code, without requiring intervention from the infrastructure team.
-Infrastructure team.
-The infrastructure team is committed to optimizing GPU kernel libraries, training frameworks, and parallel backbones.
-With XPUTimer’s lightweight logging system, the infrastructure team can gather sufficient runtime data to analyze submitted jobs and identify new optimization opportunities.
-Anomalies caused by codebase modifications that cannot be resolved by the algorithm team are ultimately routed to the infrastructure team.
-Operations team.
-The operations team is responsible for maintaining the stability of all low-level resources.
-When a training job is terminated due to low-level issues, the operations team queries intra-kernel tracing information provided by XPUTimer.
-This allows them to identify the faulty machine without requiring a fully comprehensive test.
-8.2 Holistic Diagnostics
-The training stack for advancing AIGC is inherently complex.
-Previously, in the absence of a holistic diagnostic framework like XPUTimer, significant effort was wasted on communication among algorithm, infrastructure, and operations teams.
-Algorithm teams often required collaboration with the infrastructure team to address slowdowns caused by minor but inefficient codebase modifications.
-When low-level issues arose, the operations team had to conduct extensive low-level benchmarking tests to identify the problem, resulting in delays to the algorithm teams’ training jobs.
-The holistic diagnostics provided by XPUTimer streamline this process by pinpointing elementary inefficiencies in the codebase for algorithm teams, only routing complex anomalies with aggregated metrics to the infrastructure team, and supplying the operations team with valuable runtime low-level data.
+By leveraging Flare, algorithm teams can independently identify and resolve regression anomalies caused by inefficient code without requiring intervention from the infrastructure team.
+With Flare’s lightweight logging, the infrastructure team also collects sufficient runtime data to analyze submitted jobs and independently discover new optimization opportunities.
+Only regression anomalies that cannot be resolved by algorithm teams are routed to the infrastructure team.
+Overall, the frequency of collaboration on recurrent regressions, such as unnecessary synchronizations, decreased by 63.5% within a one-week deployment, accelerating model evolution for algorithm teams and allowing the infrastructure team to focus more on system optimization.
 8.2 Using Historical Data
-Historical traces are essential for enhancing the effectiveness of anomaly detection.
-While runtime data is crucial, it is insufficient on its own for accurately identifying anomalies.
-XPUTimer detects various issues by comparing real-time data against historical data.
-For instance, when using issue latency distributions to detect kernel-issue stalls, XPUTimer relies on historical data from specific backbones operating on specific hardware.
-By collecting historical data from healthy jobs, XPUTimer can efficiently identify anomalies in newly submitted LLM training jobs.
-In the future, as deployment data grows, we aim to release relevant datasets to further streamline the diagnosis of LLM training.
-8.4 Hardware Extensibility
-Currently, XPUTimer does support the extensibility of additional hardware, particularly NPUs dedicated to DL training.
-Its tracing mechanism is not only backbone-agnostic but also designed for seamless extension to other NPUs.
-Since XPUTimer directly instruments key code segments at the Python and C++ runtime levels, extending it is straightforward, provided the relevant computation kernel is supplied by the NPU vendor.
-8.5 Diagnosing Communication Slowdown
-Diagnosing communication slowdowns caused by issues such as network jitter is challenging, as runtime tracing can introduce significant overhead or lack accuracy.
-Currently, XPUTimer utilizes a method similar to MegaScale [2], leveraging NCCL tests to identify faulty machines or switches.
-We are in the process of developing an eBPF-based [49] tracing tool to accurately monitor bandwidth across RDMA NICs with minimal overhead.
-In the future, XPUTimer will be further enhanced with these fine-grained tracing capabilities.
+Historical traces are essential for enhancing the effectiveness of regression detection.
+While runtime data is crucial, it is insufficient on its own for accurate regression diagnosis.
+Flare detects issues by comparing real-time data against historical data.
+E.g., when using issue latency distributions to detect kernel-issue stalls, Flare relies on historical data from specific backends operating on specific hardware in Figure 11.
+However, regression detection is inherently difficult to ascertain, even with historical data.
+Therefore, in practice, we adopt a conservative policy: we report potential regressions and their root causes without directly terminating training jobs.
+8.3 Hardware Extending
+Currently, Flare has supported the extensibility of additional hardware, particularly NPUs dedicated to DL training.
+Since Flare directly instruments key code segments at the Python and C++ runtime levels, extending it is straightforward.
+We already support internal CUDA-native NPUs, incurring less than $0.5\%$ runtime overhead when running on 450 NPUs.
+We also investigate the extensibility of Flare’s intra-kernel inspecting on NPUs.
+Our study shows that the methodology is largely extensible, as NPUs also use dedicated hardware cores for cross-device communication.
+Inspecting their register status provides additional information for the operations team to address hang-error anomalies, but this requires additional engineering effort.
+8.4 Scopes & Limitations
+Flare operates within certain scopes and has limitations.
+Firstly, Flare targets large-scale private GPU clusters for LLM training and relies on historical data to calibrate its diagnostic metrics.
+This assumption is reasonable.
+Accessing historical data is practical in a private GPU clusters.
+In our internal cluster, we also profiled typical LLMs and parallel backends at different scales to obtain ground-truth data for building the diagnostic mechanism.
+Secondly, Flare cannot detect regressions in training jobs with major architecture updates, e.g., switching from a dense LLM to a sparse MoE LLM.
+In such cases, Flare must collect new historical data to refine its diagnostic engine.
+This requirement is usually reasonable, as Flare is designed to relieve the infrastructure team from repetitive efforts on recurrent regression anomalies, allowing them to focus instead on optimizing LLMs with new architectures.
 9 Related Work
 Anomaly diagnosis in distributed training.
-Anomaly diagnosis in large-scale distributed deep learning training tasks has consistently been a hot research focus[2, 14, 3, 50, 51, 52].
-Megascale [2] only focuses on LLM training tasks based on Megatron-LM.
+Anomaly diagnosis in large-scale distributed deep learning training tasks has consistently been a hot research focus[jiangMegaScaleScaling, wuGREYHOUNDHunting, dongEnhancingLargescale, peng2018optimus, haider2011fault, chen2020elastic].
+Megascale [jiangMegaScaleScaling] only focuses on LLM training tasks based on Megatron-LM.
 It identifies network-related hardware and software issues in the training process by conducting intra-host network tests and NCCL tests.
-Falcon [14] detects prolonged iterations using the Bayesian Online Change-Point Detection algorithm.
-C4D [3] modifies the Collective Communication Library to collect message statistics, such as sizes and durations of transfers, to identify the performance bottlenecks.
+Greyhound [wuGREYHOUNDHunting] detects prolonged iterations using the Bayesian Online Change-Point Detection algorithm.
+C4D [dongEnhancingLargescale] modifies the Collective Communication Library to collect message statistics, such as sizes and durations of transfers, to identify the performance bottlenecks.
 However, these approaches primarily address communication-related performance issues and fail to encompass the anomalies across the LLM training stack.
-Additionally, they depend heavily on a single parallel backbone, Megatron, which limits their generality.
-Meanwhile, many researches[50, 51, 52] focus on task recovery.
-These approaches are orthogonal to XPUTimer, which could seamlessly integrate with XPUTimer.
+Additionally, they depend heavily on a single parallel backend, limiting their generality.
 Anomaly diagnosis in large-scale datacenter.
-Many research efforts[53, 54, 55, 56, 57] focus on anomaly diagnosis at different levels of the datacenter, including runtime, network, and storage.
-SCALENE[53] introduces an algorithm to assist Python programmers in optimizing their code by distinguishing between inefficient Python execution and efficient native execution.
-AND[54] is a unified application-network diagnosing system that leverages a single metric, TCP retransmissions (TCP retx), to identify network anomalies in cloud-native scenarios.
-AAsclepius[55] proposes a PathDebugging technique to trace fault linkages between the middle network and autonomous systems.
-Researchers [56] from Alibaba analyze four key factors that impact SSD failure correlations: drive models, lithography, age, and capacity.
-As these works address anomaly problems in specific scenarios, they are unable to resolve the challenges targeted by XPUTimer in large-scale distributed LLM training.
+Many research efforts[bergerTriangulatingPython, wangDiagnosingApplicationnetwork, yangAAsclepiusMonitoring, han2021depth, wangZeroOverhead] focus on anomaly diagnosis at different levels of the datacenter, including runtime, network, and storage.
+AND[wangDiagnosingApplicationnetwork] is a unified application-network diagnosing system that leverages a single metric, TCP retransmissions (TCP retx), to identify network anomalies in cloud-native scenarios.
+AAsclepius[yangAAsclepiusMonitoring] proposes a PathDebugging technique to trace fault linkages between the middle network and autonomous systems.
+Researchers [han2021depth] from Alibaba analyze four key factors that impact SSD failure correlations: drive models, lithography, age, and capacity.
+As these works address anomaly problems in specific scenarios, they are unable to resolve the challenges targeted by Flare in large-scale distributed LLM training.
 10 Conclusion
-In this paper, we introduce XPUTimer, a real-time diagnostic framework for LLM training.
-By addressing critical challenges such as lightweight long-term monitoring, root cause detection, and backbone extensibility, XPUTimer provides a comprehensive solution for anomaly diagnostics across large-scale GPU clusters.
-With its novel intra-kernel tracing mechanism and holistic aggregated metrics, XPUTimer not only reduces diagnostic complexity but also improves the detection and resolution of non-obvious slowdowns and errors.
+In this paper, we introduce Flare, a real-time diagnostic framework for LLM training.
+By addressing challenges such as lightweight long-term monitoring, root cause detection, and backend extensibility, Flare provides a comprehensive solution for anomaly diagnostics across large-scale GPU clusters.
+With its novel intra-kernel inspecting mechanism and holistic aggregated metrics, Flare not only reduces diagnostic complexity but also improves the detection and resolution of non-obvious slowdowns and errors.
 Its deployment in real-world training clusters, spanning over 6,000 GPUs, demonstrates its efficacy in diagnosing distributed LLM training.
-Furthermore, XPUTimer is open-sourced to encourage broader adoption.
+Acknowledgments
+This work is partially sponsored by the National Key Research and Development Program of China (2024YFB4505700), National Natural Science Foundation of China (62232011) and Natural Science Foundation of Shanghai Municipality (24ZR1430500) and CAAI-Ant Group Research Fund.
+We thank the anonymous reviewers and our shepherd Andrew Moore for their constructive feedback and suggestions.